### v1

# Defensive Model Expansion for Robust Bayesian Inference

Abstract
Some applied researchers hesitate to use nonparametric methods, worrying that they will lose power in small samples or overfit the data when simpler models are sufficient. We argue that at least some of these concerns are unfounded when nonparametric models are strongly shrunk towards parametric submodels. We consider expanding a parametric model with a nonparametric component $r(x)$ that is heavily shrunk toward zero. This construction allows the model to adapt automatically: if the parametric model is correct, the nonparametric component disappears, recovering parametric efficiency, while if it is misspecified, the flexible component activates to capture the missing signal. We show that this adaptive behavior follows from simple and general conditions. Specifically, we prove that Bayesian nonparametric models anchored to linear regression, including variants of Gaussian processes regression and Bayesian additive regression trees, consistently identify the correct parametric submodel when it holds and give asymptotically efficient inference for regression coefficients. In simulations, we find that the general BART model performs identically to correctly specified linear regression when the parametric model holds, and substantially outperform it when nonlinear effects are present. This suggests a practical paradigm: defensive model expansion as a safeguard against model misspecification.
1 Introduction
In our experience, non-statistician collaborators are often cautious about using nonparametric methods. Many of their reasons for skepticism are reasonable: “black-box” procedures based on machine learning can be hard to interpret and, even when they improve predictions, may not lead to clear scientific understanding or actionable policy recommendations. Another common concern is that nonparametric methods may not justify the extra computational effort or expertise they require, relative to their impact on final conclusions.
In our view, however, some criticisms of nonparametric methods are less well-founded. One persistent worry is that nonparametric methods lose power in small samples, making it risky to use them out of fear that a “statistically significant” result might disappear. This need not be true: to the extent that they reduce error variability, nonparametric methods can actually increase power. And even if some power were lost, one could argue that including a nonparametric component provides a more honest reflection of modeling uncertainty. Another frequent complaint is that nonparametric models tend to overfit, and that parsimonious parametric models are safer. While overfitting can occur with poorly regularized nonparametric models, it is not an inherent problem.
Our goal is to show that properly constructed Bayesian nonparametric methods can perform just as well as parametric methods, even in small samples with high signal-to-noise ratios, by using a strategy of defensive model expansion. The main idea is to add a flexible adjustment term $r(x)$, with shrinkage toward zero, such that when $r(x)\equiv 0$ we recover the original parametric model. This approach has several advantages:
  • By incorporating model selection into the modeling process itself, we avoid relying on an analyst’s ad hoc (and often data-dependent) choices about which interactions or nonlinearities to include. This reduces researcher degrees of freedom and guards against garden of forking paths problems at the analysis stage (Gelman and Loken,, 2013).
  • If the parametric model is badly misspecified or omits important signals, expanding into a nonparametric model can improve power. With a flexible “alternative” model where $r(x)\neq 0$, we can also use the estimated size of $r(\cdot)$ as a diagnostic for parametric model misspecification: if $|r(x)|$ is large on average then this is evidence that the parametric submodel model fits poorly.
  • If the nonparametric component turns out to be important, we obtain more honest uncertainty quantification and the chance for richer scientific discoveries. Moreover, if we specify a fully Bayesian model with sensible priors in advance, then we have still only “used the data once” in performing inference on these discoveries (Woody et al.,, 2021).
  • When the parametric model is correct, we lose little power. In fact, we will show that we can still obtain asymptotically efficient parameter estimation.
The downsides of this strategy are minimal, as we can always shrink the nonparametric component heavily toward the parametric submodel. Often the result will simply be to confirm that the nested parametric model is a good approximation to reality. While this may seem anti-climactic, it is still useful information. Occasionally, the strategy may be too deferential to the parametric submodel, but still detect obvious signs of misspecification; we illustrate this in Section 4, where correctly specified parametric models achieve predictive performance that is matched by nonparametric models. In large samples, a common outcome is that we identify interesting ways to expand the nested parametric model in a Bayes-valid fashion, which we see in Section 4.4.
1.1 Existing Literature
It has long been recognized that one can design Bayesian models that adaptively become more complex when the data demand it. This idea is a core theme in the Bayesian nonparametric literature, although in practice much of the focus has been on clustering and latent feature models (Orbanz and Teh,, 2011; Gershman and Blei,, 2012; Ghahramani et al.,, 2007). A key difference from the settings we consider is that clustering and latent feature problems are evidently complex (with the main question being only how complex), whereas we focus on settings that are plausibly simple, where a parametric model might in fact be sufficient. See Li and Dunson, (2020) for an example of using a baseline nonparametric model as a tool for evaluating parametric models, as we do here.
We briefly describe some other works that have been particularly influential on our approach. Neal, (1995) makes the point that, “There is no provision in the Bayesian framework for changing the model or the prior depending on how much data was collected. If the model and prior are correct for a thousand observations, they are correct for ten observations as well (though the impact of using an incorrect prior might be more serious with fewer observations),” and that when the underlying phenomena is evidently complex the appropriate solution to finding a simpler model outperforming a more complex model is “to design a different complex model that captures whatever aspect of the problem led to the simple model performing well.”
In the context of mixture models, a selling point of infinite mixtures $f(y)=\sum_{k=1}^{\infty}\omega_{k}\,g(y\mid\vartheta_{k})$ is that they can collapse to a single parametric model $g(y\mid\theta)$ when most prior mass is placed on the event $\omega_{1}\approx 1$. For Dirichlet process mixtures (Escobar and West,, 1995), this is accomplished by shrinking the concentration parameter heavily towards zero. For conditional mixture models — such as dependent Dirichlet processes (DDPs, MacEachern,, 2000; Quintana et al.,, 2022) — this allows us to recover standard parametric generalized linear models when appropriate (Shahbaba and Neal,, 2009; Hannah et al.,, 2011).
Our strategy also represents a type of continuous model expansion in the sense of Gelman and Shalizi, (2013), who recommend “forming a larger model that includes both A and B as special cases,” as an alternative to testing Model A against Model B. Like them, we do not formally test the parametric submodel (except as a theoretical device). While they explicitly do not recommend Bayesian nonparametrics as a tool for continuous model expansion, we also do not feel that our proposal is at odds with the additional use of model checking they recommend, and do not view our use of an additional nonparametric component to be a complete solution to model misspecification. Related works include Kamary et al., (2018) and Yao et al., (2018).
1.2 Contributions
We recommend defensive model expansion partly because, under very general conditions, it yields efficient inference when the associated parametric model is correctly specified. The story is particularly simple if we replace the posterior distribution with a fractional posterior (Walker and Hjort,, 2001; O’Hagan,, 1995), which allows us to sidestep the technical testing and entropy conditions required by Ghosal et al., (2000). In fact, the nonparametric component of the prior plays little direct role in our analysis, making the results extremely general.
We begin by presenting general results on posterior concentration for nonparametric models when parametric submodels have sufficient prior mass. With the exception of Theorems 6 and 7, these are straightforward extensions of existing results for fractional posteriors. We prove results for convergence rates, model selection, and asymptotic normality under the assumption that the parametric model is correctly specified. Our model selection results essentially show that (i) we can correctly select the parametric model and (ii) the posterior distribution satisfies a Bernstein–von Mises theorem. We also prove two semiparametric Bernstein–von Mises theorems that apply in settings where model selection consistency is difficult to use, which guarantees efficient posterior concentration of a projection parameter $\theta^{\star}$ around its true value. Such projection parameters are often used to summarize posterior distributions (Woody et al.,, 2021).
We then apply these results to Gaussian process (GP) and Bayesian additive regression tree (BART) models. We show that BART models naturally satisfy the conditions required for a Bernstein–von Mises result (see also Rocková,, 2020), as do “spike-and-GP” priors. While we do not study this here, our semiparametric Bernstein–von Mises results for GPs also have scope for being extended to projection parameters even when the anchoring parametric model is misspecified. These theoretical results are validated on simulated data.
In addition to being theoretically simple, defensive model expansion is easy to implement in practice. For example, because our convergence rate and model selection results apply directly to BART models, existing software for the general BART model of Tan and Roy, (2019) makes it straight-forward to obtain parametric adaptivity with default prior specifications. BART models are especially attractive because, in addition to adapting to parametric structures, they also adjust to locally varying smoothness and the presence of low-order interactions (Linero and Yang,, 2018; Ročková and van der Pas,, 2020; Jeong and Rockova,, 2023).
1.3 Notation
We let $Z_{1},Z_{2},\ldots$ denote a sequence of random variables taking values in a set $\mathcal{Z}\subseteq\mathbb{R}^{Q}$ and write $\boldsymbol{Z}_{N}=(Z_{1},\ldots,Z_{N})$ for the data. We consider a model $\mathcal{F}=\{F_{\theta}:\theta\in\Theta\}$ for some parameter space $\Theta$ such that the $F_{\theta}$’s are mutually dominated by some measure $dz$ with associated densities $f_{\theta}(z)$. We assume that $Z_{i}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}f_{\theta_{0}}$ for some $\theta_{0}\in\Theta$, although we will also briefly discuss the setting where the model is misspecified. For any function $g$, let $\mathbb{E}_{\theta}\{g(Z)\}$ and $\operatorname{Var}_{\theta}\{g(Z)\}$ denote expectation and variance under $f_{\theta}$.
We assume that $\theta$ decomposes as $(r,\eta)$ where $r\in\mathcal{H}$ is infinite-dimensional and $\eta\in E\subseteq\mathbb{R}^{P}$ is-finite dimensional, such that $\Theta=\mathcal{H}\times E$. The family $\mathcal{F}_{0}=\{F_{\theta}:\eta\in E\ \text{and}\ r=r_{0}\}$ represents a parametric submodel of $\mathcal{F}$. In our examples, $r(\cdot)$ is a mapping $\mathbb{R}^{P}\to\mathbb{R}$ and $r_{0}(\cdot)\equiv 0$. Intuitively, $\eta$ corresponds to the usual parametric component (e.g., regression coefficients), while $r$ is a flexible nonparametric adjustment. We let $\theta_{0}=(r_{0},\eta_{0})$.
We let $\Pi(\cdot)$ denote a prior on $\Theta$, $L(\theta)=\prod_{i}f_{\theta}(Z_{i})$ the likelihood function, and $R(\theta)=L(\theta)/L(\theta_{0})$ the likelihood ratio. The (fractional) posterior is defined by $\Pi_{\alpha}(A\mid\boldsymbol{Z}_{N})=\int_{A}L(\theta)^{\alpha}\ \Pi(d\theta)/\int L(\theta)^{\alpha}\ \Pi(d\theta)$ for $\alpha\in(0,1]$.
Product Prior Assumption (PPA)
Throughout, we mostly assume independent priors on $r$ and $\eta$, i.e., $\Pi=\Pi_{r}\times\Pi_{\eta}$, with $\Pi_{\eta}$ admitting a continuous density $\pi_{\eta}$ on $E$ such that $\pi_{\eta}(\eta_{0})>0$. We also assume that $\mathcal{H}$ is a Hilbert space with inner product $\langle\cdot,\cdot\rangle$, and that there exists a norm $\|\cdot\|_{\mathcal{H}}$ and positive constants $(C(\theta_{0}),\epsilon)$ such that $\max\{K(\theta_{0}\|\theta),V(\theta_{0}\|\theta)\}\leq C(\theta_{0})\{\|\eta-\eta_{0}\|^{2}+\|r-r_{0}\|_{\mathcal{H}}^{2}\}$ for all $\theta\in\Theta$ satisfying $\|\eta-\eta_{0}\|^{2}+\|r-r_{0}\|_{\mathcal{H}}^{2}\leq\epsilon^{2}/C(\theta_{0})$. Here $K(\theta_{0}\|\theta)$ and $V(\theta_{0}\|\theta)$ are Kullback–Leibler type divergences, as defined in Section 3.
We now describe three concrete models where our results apply. We will focus mainly on Model NR for clarity.
Nonparametric Regression (Model NR)
Let $Z_{i}=(X_{i},Y_{i})$ with $Y_{i}\mid X_{i}=x]\sim\operatorname{Normal}\{x^{\top}\beta_{0}+r_{0}(x),\sigma_{0}^{2}\},\text{ }X_{i}\sim F_{X},$ where $\theta=(r,\beta,\sigma)$ and $X_{i}\in[0,1]^{D}$ for some $D$. Anchoring to the parametric submodel with $r_{0}(x)\equiv 0$ recovers ordinary linear regression. Our goals are to estimate the regression function $\mu_{0}(x)=x^{\top}\beta_{0}$ and, ideally, recover $\beta_{0}$ efficiently.
Nonparametric Classification (Model NC)
We adopt the same framework as Model NR, except $Y_{i}\in\{0,1\}$ and $\mathbb{E}_{\theta}(Y_{i}\mid X_{i}=x)=\mu(x)=(1+\exp\{x^{\top}\beta+r(x)\})^{-1}$ so that $Y_{i}\mid X_{i}=x]$ is a logistic regression and $\theta=(r,\beta)$. Our goal again is to estimate $\mu_{0}(x)$ and estimate $\beta_{0}$ efficiently.
Density Regression (Model DR)
Similar to Model NR, assume $Y_{i}\mid X_{i}=x]\sim\operatorname{Normal}(x^{\top}\beta_{0},\sigma_{0}^{2})$ with $\theta=(r,\beta,\sigma)$, but rather than using $r(x)$ to adjust the regression function we set $f_{\theta}(y\mid x)=\frac{\Phi\{r(x,y)\}\phi(y\mid x^{\top}\beta,\sigma)}{\int\Phi\{r(x,y)\}\phi(y\mid x^{\top}\beta,\sigma)\ dy}$ where $\phi(y\mid\mu,\sigma)$ is the $\operatorname{Normal}(\mu,\sigma^{2})$ density and $\Phi(\cdot)$ is the $\operatorname{Normal}(0,1)$ distribution function. Once again, $\theta_{0}=(r_{0},\beta_{0},\sigma_{0})$ with $r_{0}\equiv 0$. When $r(x,y)$ is a Gaussian process, this model is a conditional version of the Gaussian process density sample (Murray et al.,, 2008), which is an extremely flexible model for a conditional density. It was shown by Li et al., (2023) that when $r(\cdot)$ is given a smooth BART (SBART) prior, the posterior concentrates at a near-minimax rate about $f_{\theta_{0}}$ if $r_{0}(x,y)$ is Hölder-smooth and other assumptions are satisfied. Linero et al., (2022) use this strategy in the context of Bayesian nonparametric survival analysis. Complementing the adaptive concentration results provided in those works, the results here also establish that these models can adapt to the underlying parametric regression model being correctly specified.
2 Modeling Strategies
We now describe strategies for modeling the nonparametric component $r(\cdot)$ and for assessing how much the nested parametric submodel may be misspecified. We focus on two widely used classes of Bayesian models: Bayesian additive regression trees (BART) and Gaussian processes (GPs).
2.1 General Bayesian Additive Regression Trees Models
For routine use, we recommend the Bayesian Additive Regression Trees (BART, Chipman et al.,, 2010) prior for the nonparametric component $r(\cdot)$. In the context of Model NR the regression function is $\mu(x)=x^{\top}\beta+\sum_{t=1}^{T}g(x;\mathcal{T}_{t},\mathcal{M}_{t})$ where each $g(x;\mathcal{T}_{t},\mathcal{M}_{t})$ is a regression tree, $\mathcal{T}_{t}$ denotes the tree structure, and $\mathcal{M}_{t}=\{\lambda_{\ell t}:\ell=1,\ldots,L_{t}\}$ is the set of leaf node predictions. Tan and Roy, (2019) refer to this as a general BART (GBART) model, and it can be fit using the SoftBart package (Linero,, 2022) using the gsoftbart_regression function. A schematic of the construction of the model is given in Figure 1, which shows a single tree from the ensemble and the associated step function.
Figure 1: Left: an example of a regression tree with input $x=(x_{1},x_{2})$ supported on $[0,1]$. Right: the induced step function on $0,1]^{2}$ .
We now review the prior specification of BART, which is important for our purposes. The default recommendation of Chipman et al., (2010) takes $\lambda_{t\ell}\sim\operatorname{Normal}(0,\sigma^{2}_{\mu}/T)$ with $\sigma_{\mu}=0.5$ after centering and scaling the $Y_{i}$’s to lie between -0.5 and 0.5. The scaling of $\sigma^{2}_{\mu}$ by $T$ is done so that $\operatorname{Var}\{\sum_{t}g(x;\mathcal{T}_{t},\mathcal{M}_{t})\}=\sigma^{2}_{\mu}$, which in turn leads to a Gaussian process limit as $T\to\infty$ (Linero,, 2017). One way to shrink the GBART model to a linear regression is to place a prior on $\sigma_{\mu}$ supported near 0, such as a half-normal or half-Cauchy prior.
Another way for GBART to reduce to a linear regression is for all of the trees in the ensemble to be “empty” in the sense of not splitting on any variables. In Section 3.2 we will show that this suffices for GBART to possess strong theoretical properties when the linear model is correct. Chipman et al., (2010) propose a branching process prior for $\mathcal{T}_{t}$ that (i) decides whether each node at a given depth of the tree will be converted into a branch node with two leaf children; (ii) iterating over the depths; and (iii) terminating if all of the nodes at a given depth are leaf nodes. The probability of a node at depth $d$ being a branch node under the prior is specified to be $q(d)=a/(1+b)^{d}$ where by default $a=0.95$ and $b=2$. In particular, the probability that a tree consists of no splits is (1a)(1-a), and the prior probability that the GBART model reduces to the linear model is therefore $(1-a)^{T}>0$.
2.2 Gaussian Process Regression
A common approach for anchoring Model NR to a parametric submodel is to use the Gaussian process regression (GP regression) model. GPs are distributions on random functions, and can be used as priors model for regression functions $r(\cdot)$. A GP is determined by a mean function $m(x)$ and a positive semi-definite kernel $\kappa(x,x^{\prime})$, with $\mathbb{E}_{\Pi}r(x)=m(x)$ and $\operatorname{Cov}_{\Pi}\{r(x),r(x^{\prime})\}=\kappa(x,x^{\prime})$. We write $\operatorname{GP}(m,\kappa)$ for the associated Gaussian process. For a review of Gaussian processes and alternative choices of $\kappa(\cdot,\cdot)$, see Rasmussen and Williams, (2006). Importantly, GPs are conjugate priors for $\mu(x)$ in the model $Y_{i}\mid X_{i}=x]\sim\operatorname{Normal}\{\mu(x),\sigma^{2}\}$ when $\kappa(x,x^{\prime})$ is held fixed.
For model NR, we take $r\sim\operatorname{GP}\{0,\sigma^{2}_{\mu}\,\kappa(\cdot,\cdot)\}$. Equivalently, letting $\mu(x)=x^{\top}\beta+r(x)$ we have $\mu\sim\operatorname{GP}(m,\sigma^{2}_{\mu}\,\kappa)$ where $m(x)=x^{\top}\beta$. If $\beta\sim\operatorname{Normal}(0,\sigma^{2}_{\beta})$ this simplifies further to $\mu\sim\operatorname{GP}(0,\sigma^{2}_{\mu}\kappa+\sigma^{2}_{\beta}\kappa_{\text{linear}})$ where $\kappa_{\text{linear}}(x,x^{\prime})=x^{\top}x^{\prime}$. In our illustrations we use the squared exponential function $\kappa_{\text{SE}}(x,x^{\prime})=\exp(-\rho\|x-x^{\prime}\|^{2})$, with the lengthscale $\rho$ controlling the smoothness of samples from the prior.
Just as with BART, there are multiple ways to obtain adaptation to parametric structures. Below, we describe some strategies for ensuring that a GP is capable of attaining adaptivity to $r_{0}(x)=0$.
Scale Mixture Models
One option is to place a hyperprior on $\sigma_{\mu}$, with heavy shrinkage towards zero. We call this the scale mixture of GPs approach. Intuitively, this lets the model collapse to the parametric submodel when the nonparametric adjustment is unnecessary by taking $\sigma_{\mu}\approx 0$. When performing inference on projection parameters (see Section 2.4) this strategy is sufficient for most purposes.
Spike and Slab Models
A second option is to use a “spike-and-slab” prior (George and McCulloch,, 1993) with a GP slab distribution:
$r\sim p_{0}\,\delta_{r_{0}}+(1-p_{0})\,\operatorname{GP}(m,\kappa),$
where $\delta_{r_{0}}$ denotes a point mass at $r_{0}(x)\equiv 0$ and $0<p_{0}<1$. We refer to this as the spike-and-GP approach. If $p_{0}$ is large then the model strongly favors the parametric submodel, while still allowing flexibility if supported by the data. In Section 3.1 we will see that spike-and-slab priors adapt to parametric structures with essentially no further requirements.
Semiparametric Bernstein-von Mises for Projections
A third strategy is to simply use a GP for $r(x)$ and then perform inference on a projection parameter $\eta^{\star}$ obtained from projecting the full model onto the parametric submodel $\mathcal{F}_{0}$. Concretely, we might compute the Kullback-Leibler projection $\theta^{\star}=\arg\min_{\theta\in\mathcal{F}_{0}}K(\theta\|\theta^{\star})$ for each sample $\theta$ from the posterior and, because $\theta_{0}\in\mathcal{F}_{0}$, hopefully $\theta^{\star}$ will concentrate faster around $\theta_{0}$ than $\theta$. This can be used with either of the previous two approaches, but sometimes works on its own as well. When the parametric model is correctly specified, appropriately chosen GP priors can yield efficient inference. However, poorly chosen priors may still lead to inefficient or invalid inference. For example, we will see that even if the posterior converges at a near-parametric rate such as $N^{-1/2}(\log N)^{t}$ for some $t>0$, the bias in estimating $\beta_{0}$ may decay very slowly, making inference inefficient in finite samples. This occurs, for example, when we incorrectly set $\beta=0$ and model $r(x)$ with a squared exponential kernel.
2.3 Orthogonalization
When using models with predictions of the form $X_{i}^{\top}\beta+r(X_{i})$, such as Model NR or Model NC, it is helpful to require the nonparametric component $r(x)$ to be orthogonal to the linear predictors. That is, we require that $\sum_{i}r(X_{i})\,X_{i}=\mathbf{0}$. This condition serves two purposes. First, it can substantially improve the mixing of MCMC algorithms used to fit the model. Second, it prevents $r(\cdot)$ from becoming collinear with $X_{i}^{\top}\beta$, making it easier to shrink $r(\cdot)$ toward zero when it is unnecessary. Orthogonalization is not required for our theoretical results to hold, but it is often useful in practice. For GP models, orthogonalization is implemented by replacing the kernel $\kappa(x,x^{\prime})$ with the projected kernel $\kappa^{\star}(x,x^{\prime})=\kappa(x,x^{\prime})-k_{x}^{\top}\boldsymbol{X}(\boldsymbol{X}^{\top}\boldsymbol{K}\boldsymbol{X})^{-1}\boldsymbol{X}^{\top}k_{x^{\prime}},$ where $\boldsymbol{K}_{ij}=\kappa(X_{i},X_{j})$, $k_{x}=(\kappa(x,X_{1}),\ldots,\kappa(x,X_{N}))^{\top}$, and $\boldsymbol{X}\in\mathbb{R}^{N\times P}$ is the design matrix (Plumlee and Joseph,, 2018).
2.4 Summarizing Nonparametric Fits
Our proposed workflow for using these models, illustrated in Section 4.4, is to:
  1. 1.
    Obtain posterior samples from the larger nonparametric model.
  2. 2.
    Project these samples onto the parametric submodel $\mathcal{F}_{0}$ and assess its adequacy.
  3. 3.
    If the parametric model is inadequate with high posterior probability, fit additional summaries (such as additive models or tree-based fits) to identify specific aspects of the misspecification, and use these to refine $\mathcal{F}_{0}$.
One practical implementation is to form a linear projection summary by minimizing
$\text{SSE}=\sum_{i=1}^{N}\{\widehat{\mu}(X_{i})-X_{i}^{\top}\beta^{\star}\}^{2},$ (1)
where $\widehat{\mu}(x)$ is the posterior mean function (or a posterior draw). We then compute $R^{2}=1-\frac{\text{SSE}}{\sum_{i}\{\widehat{\mu}(X_{i})-\bar{\mu}\}^{2}},\text{ where }\bar{\mu}=N^{-1}\sum_{i}\mu(X_{i}).$ If $R^{2}$ is sufficiently close to 1, we conclude that the parametric model is adequate. The required closeness depends on context: if prediction is the main goal, we might demand $R^{2}$ very close to 1, while if the goal is only a broad scientific conclusion, a more modest $R^{2}$ could suffice. If the linear projection is inadequate, we expand to an additive model $\mu(x)\approx\sum_{j}\gamma_{j}(x_{j})$ and, if that is still insufficient, apply a CART fit to residuals to detect missing interaction terms. For logistic regression, it is natural instead to minimize the Kullback–Leibler divergence between the fitted probabilities $\mu(X_{i})$ and a logistic regression model with predictor $X_{i}^{\top}\beta^{\star}$.
Importantly, we do not recommend formal hypothesis testing of whether $r=r_{0}$. First, we may lack power to detect $r=r_{0}$ even if the model technically allows for it. Second, even when $r\neq r_{0}$ with high posterior probability, the parametric submodel may still be practically indistinguishable from the nonparametric model in terms of its predictions.
A criticism of the above strategy is that it is not specific enough, requiring subjective judgments about whether a summary is “adequate” or not. On the contrary, we feel that this is a strength, as it forces the analyst to engage with the question of what purpose the fitted model is supposed to serve and whether it is sufficient for that purpose. Judgments, such as what an appropriate tradeoff between interpretability and predictive accuracy are, should chiefly be determined in collaboration with subject matter experts.
3 Theoretical Results
We now establish some basic theoretical properties for the BART and GP models. To formalize the idea that a properly anchored Bayesian nonparametric model can perform “just as well” as a parametric submodel we will use arguments based on the rate of convergence of a posterior distribution; the rate will either be used as a justification on its own, or as part of the assumptions for stronger justifications. We will say that the posterior convergence rate is (upper bounded by) $\epsilon_{N}$ if we have $\Pi_{\alpha}\{H(\theta_{0},\theta)>M_{N}\,\epsilon_{N}\mid\boldsymbol{Z}_{N}\}\to 0$ in probability for all sequences $M_{N}\uparrow\infty$ (where $M_{N}$ is allowed to diverge arbitrarily slowly). Here, $H(\theta_{0},\theta)=[\int\{\sqrt{f_{\theta}(z)}-\sqrt{f_{\theta_{0}}(z)}\}^{2}\ dz]^{1/2}$ denotes the Hellinger distance between $f_{\theta}$ and $f_{\theta_{0}}$. For example, we will show that both GBART and the spike-and-GP approaches give $\epsilon_{N}=N^{-1/2}$ rates and that posterior inferences for $\beta_{0}$ are efficient, provided that the parametric submodel is correct.
The most important condition governing whether parametric adaptivity is possible is whether or not the prior assigns enough mass to neighborhoods of $\theta_{0}$. In nonparametric setups, the relevant neighborhoods are defined in terms of the Kullback-Leibler (KL) neighborhoods:
$\theta\in B_{\epsilon}\quad\text{if}\quad K(\theta_{0}\|\theta)\stackrel{{\scriptstyle\text{def}}}{{=}}\mathbb{E}_{\theta_{0}}\log\frac{f_{\theta_{0}}(Z)}{f_{\theta}(Z)}\leq\epsilon^{2}\quad\text{and}\quad V(\theta_{0}\|\theta)\stackrel{{\scriptstyle\text{def}}}{{=}}\operatorname{Var}_{\theta_{0}}\log\frac{f_{\theta_{0}}(Z)}{f_{\theta}(Z)}\leq\epsilon^{2}.$
We expect to be able to estimate $\theta_{0}$ at a rate $\epsilon_{N}$ if the following prior thickness condition holds:
Definition 1 ($\epsilon_{N}$-thickness).
The prior $\Pi$ is said to be $\epsilon_{N}$-thick at $\theta_{0}$ if there exist $C_{1},C_{2}>0$ such that
$\Pi(B_{\epsilon_{N}})\geq C_{1}\,\exp(-C_{2}N\epsilon_{N}^{2}).$
We simplify our arguments by restricting attention to the $\alpha$-fractional posterior given by
$\Pi_{\alpha}(d\theta\mid\boldsymbol{Z}_{N})=\frac{L_{N}(\theta)^{\alpha}\ \Pi(d\theta)}{\int L_{N}(\theta)^{\alpha}\ \Pi(d\theta)},\qquad\alpha\in(0,1),$ (2)
where $L_{N}(\theta)=\prod_{i=1}^{N}f_{\theta}(Z_{i})$ is the likelihood. Works that have studied the properties of fractional posteriors include O’Hagan, (1995); Walker and Hjort, (2001); Bhattacharya et al., (2019); Martin and Walker, (2019). Remarkably, under essentially no further assumptions, the prior thickness condition is sufficient to guarantee posterior concentration at the rate $\epsilon_{N}$. Section 2.2 of Castillo, (2024) is particularly instructive for understanding why these conditions are sufficient and how the arguments can be extended to non-iid settings.
Interpretation of $\alpha$
The role of $\alpha$ can be understood in two ways. First, it is as if we “hold out” $100(1-\alpha)\%$ of the data, which makes the posterior less prone to overfitting (O’Hagan,, 1995). Second, as noted by Walker and Hjort, (2001), the fractional posterior is equivalent to using the full likelihood $L_{N}(\theta)$ with a data-adaptive prior $\Pi^{\star}(d\theta)\propto\Pi(d\theta)/L_{N}(\theta)^{1-\alpha}$. This construction downweights parameter values that fit the observed data too well, again limiting overfitting. This robustness to overfitting is what allows us to avoid conditions on the size of the effective support required of the full posterior.
Identifiability
Prior thickness ensures convergence in Hellinger distance, but does not automatically imply efficient estimation of $\eta_{0}$, since $r(\cdot)$ may be confounded with $\eta$. For example, in Models NR and NC, $\mu(\cdot)\to\mu_{0}(\cdot)$ does not necessarily imply $\beta\to\beta_{0}$. One way to ensure recovery of $\eta_{0}$ is through a metric identifiability condition such as $\|\eta-\eta_{0}\|\lesssim H(\theta,\theta_{0}),$ which forces $H(\theta,\theta_{0})\to 0$ only if $\eta\to\eta_{0}$ at the same rate. In Model NR this can be obtained by imposing $\mathbb{E}_{\theta}\{X_{i}r(X_{i})\}=\mathbf{0}$ and assuming $\mathbb{E}(X_{i}\,X_{i}^{\top})=\Sigma_{X}$ is full-rank. Alternatively, consistent selection of the event $r=r_{0}]$ also guarantees efficient inference for $\eta_{0}$.
Non-iid Data and Finite Sample Bounds
It is possible to extend our results to allow for non-iid data by instead defining $B_{\epsilon}$ in terms of $\mathbb{E}_{\theta_{0}}\left(\log\frac{f(Z_{1},\ldots,Z_{N}\mid\theta_{0})}{f(Z_{1},\ldots,Z_{N}\mid\theta)}\right)^{k}\leq N\,\epsilon_{N}^{2}$ for $k\in\{1,2\}$, and similarly redefining the Hellinger distance. It is also possible to convert the convergence rate arguments into non-asymptotic arguments. Finite sample results are given, for example, by Bhattacharya et al., (2019).
Non-Fractional Posteriors
The usual posterior ($\alpha=1$) requires additional technical work that we feel distracts from the message of this paper. In this case, prior thickness alone is not enough; one must also control the complexity of the support of the prior, often by placing high mass on a low-entropy sieve (Ghosal et al.,, 2000). Prior support conditions serve to stop overfitting to the data. While this strategy works, it places much stronger restrictions on the choice of $\Pi_{r}(dr)$ and can introduce an extra $(\log N)^{t}$ factor in the rates that can be cumbersome to remove.
3.1 General Results
We start with general theoretical results that will later be applied to BART and GP regression models. Throughout this section (and Sections 3.2 and 3.3) we fix $\alpha\in(0,1)$. Our first result is a general fractional posterior concentration theorem, implied for example by Theorem 2.1 of Castillo, (2024).
Theorem 1 (General Rates).
Suppose that PPA holds and that for all sequences $M_{N}\uparrow\infty$ there exist positive constants $C_{1},C_{2},N^{\star}$ such that $\Pi_{r}(\|r-r_{0}\|_{\mathcal{H}}\leq M_{N}N^{-1/2})\geq C_{1}\,e^{-C_{2}\,M_{N}^{2}}$ for $N\geq N^{\star}$. Then for all sequences $M_{N}\uparrow\infty$ we have $\mathbb{E}_{\theta_{0}}\Pi_{\alpha}\{H(\theta_{0},\theta)>M_{N}\,N^{-1/2}\mid\boldsymbol{Z}_{N}\}\to 0$ as $N\to\infty$.
It is easy to design priors that satisfy the conditions of Theorem 1. In fact, any prior that assigns positive mass to the event $r=r_{0}]$ will suffice. Spike-and-slab priors (George and McCulloch,, 1993) of the form $\Pi_{r}(dr)=p_{0}\,\delta_{r_{0}}(dr)+(1-p_{0})\,\Pi^{\star}_{r}(dr)$ (with $\delta_{r_{0}}$ denoting a point mass distribution at $r_{0}$) are natural choices.
Corollary 1 (Rate for Spike-and-Slab).
Suppose that PPA holds and that $\Pi_{r}(r=r_{0})>0$. Then for all sequences $M_{N}\uparrow\infty$ we have $\mathbb{E}_{\theta_{0}}\Pi_{\alpha}\{H(\theta_{0},\theta)\geq M_{N}\,N^{-1/2}\mid\boldsymbol{Z}_{N}\}\to 0$ as $N\to\infty$.
Anchoring to the parametric model also enables consistent model selection. If the parametric submodel holds and the nonparametric part cannot achieve the parametric rate on its own, then the posterior concentrates all its mass on the parametric model.
Theorem 2 (General Model Selection).
Suppose that PPA holds and that $\Pi_{r}(dr)=p_{0}\delta_{r_{0}}(dr)+(1-p_{0})\,\Pi_{r}^{\star}(dr)$ with $0<p_{0}<1$. Define
$\Pi^{\star}_{\alpha}(A\mid\boldsymbol{Z}_{N})=\frac{\int_{A}L(\theta)^{\alpha}\ \Pi_{r}^{\star}(dr)\ \Pi_{\eta}(d\eta)}{\int L(\theta)^{\alpha}\ \Pi_{r}^{\star}(dr)\ \Pi_{\eta}(d\eta)},$
and suppose that $\mathbb{E}_{\theta_{0}}\Pi^{\star}_{\alpha}\{H(\theta_{0},\theta)<M_{N}\,N^{-1/2}\mid\boldsymbol{Z}_{N}\}\to 0$ as $N\to\infty$. Then $\mathbb{E}_{\theta_{0}}\Pi_{\alpha}(r=r_{0}\mid\boldsymbol{Z}_{N})\to 1$ as $N\to\infty$.
Remark 1.
The assumption that $\Pi^{\star}_{\alpha}(\cdot\mid\boldsymbol{Z}_{N})$ fails to achieve the parametric rate is stronger than strictly necessary, as fractional Bayes factors are typically consistent for nested parametric model selection (O’Hagan,, 1995). This assumption is not very restrictive, however, because truly nonparametric models rarely achieve parametric rates unless they place positive mass on a correctly specified submodel.
When Theorem 2 is combined with a parametric submodel $\mathcal{F}_{0}$ that is differentiable in quadratic mean (see Section 5.5 of van der Vaart,, 1998), we also obtain a Bernstein–von Mises result. The following theorem formalizes this under standard regularity conditions.
Theorem 3 (Bernstein-von Mises).
Suppose that the conditions of Theorem 2 hold and that Assumption BVM in the Supplementary Material holds. Let $\Pi_{\alpha}(d\eta\mid\boldsymbol{Z}_{N})$ denote the marginal distribution of $\eta$ under $\Pi_{\alpha}(d\theta\mid\boldsymbol{Z}_{N})$. Then
$\|\Pi_{\alpha}(d\eta\mid\boldsymbol{Z}_{N})-\operatorname{Normal}(d\eta\mid\widehat{\eta},(\alpha N)^{-1}\mathcal{I}(\eta_{0})^{-1})\|_{TV}\to 0\quad\text{in $F_{\theta_{0}}$-probability},$
where $\mathcal{I}(\eta)$ denotes the unit Fisher information in the parametric model $\mathcal{F}_{0}$, $\|\cdot\|_{TV}$ denotes total variation distance, and $\widehat{\eta}$ is any estimator such that $\sqrt{N}(\widehat{\eta}-\eta_{0})\to\operatorname{Normal}(0,\mathcal{I}(\eta_{0})^{-1})$ in distribution.
The value of Theorem 3 relative to the usual Bernstein-von Mises theorem is that it represents an oracle property of the fractional posterior: if the parametric model is correctly specified, our nonparametric prior will nevertheless obtain efficient inference for $\eta_{0}$. On the other hand, if the parametric model is misspecified we will also be able to obtain efficient nonparametric estimation rates for $\theta_{0}$; the exact rates we get under misspecification will depend on the choice of $\Pi^{\star}(dr)$ and on $r_{0}(x)$.
3.2 Results for Bayesian Additive Regression Trees
As illustrated by Ročková and van der Pas, (2020); Jeong and Rockova, (2023); Linero and Yang, (2018), BART models are adaptive to a wide variety of structures. In Model NR, for example, they are adaptive to sparsity, (locally-varying) smoothness, and low-order interaction structure in $r_{0}(x)$ out-of-the-box. Adding to this list of desirable properties, appropriately constructed BART models are also adaptive to parametric structures with no additional modifications.
Theorem 4 (BART Adaptivity).
Suppose that Model NR holds with $r_{0}(x)=0$ for all $x\in[0,1]^{D}$ and suppose that the prior $\Pi(d\theta)$ is such that
  • $(\beta,\sigma)$ has a continuous density $\pi_{\beta,\sigma}$ with $\pi_{\beta,\sigma}(\beta_{0},\sigma_{0})>0$.
  • $r$ is independent of $(\beta,\sigma)$ and has a BART prior as described in Section 2.1.
Then the conditions of Theorem 1 hold. Additionally, let $\mathcal{N}$ be the event that all decision trees in the ensemble contain no splitting rules. Then $\Pi_{\alpha}(\mathcal{N}\mid\boldsymbol{Z}_{N})\to 1$ in $F_{\theta_{0}}$-probability.
This result is close in spirit to Theorem 2. While we cannot directly apply Theorem 3, Theorem 4 is strong enough to imply a Bernstein–von Mises result.
Theorem 5 (BART Bernstein-von Mises).
Suppose that the conditions of Theorem 4 hold and that $F_{X}$ is such that $X_{i1}\equiv 1$ and that $\mathbb{E}(X_{i}\,X_{i}^{\top})=\Sigma_{X}$ is non-singular. Define the vector $\eta=(\beta_{1}+r(0),\beta_{2},\beta_{3},\ldots,\beta_{D},\sigma)^{\top}$ and let $\widehat{\eta}$ by the maximum likelihood estimator of $(\beta,\sigma)$ in Model NR when $r_{0}\equiv 0$. Then the conclusion of Theorem 3 holds.
Remark 2.
Versions of the above results can also be derived for Model NC and Model DR; we only give results for Model NR for brevity.
3.3 Results for Gaussian Processes
We have discussed two distinct strategies for Gaussian process models. The first is to use a spike-and-slab prior $\Pi(dr)=p_{0}\,\delta_{r_{0}}+(1-p_{0})\,\Pi^{\star}_{r}(dr)$ with $\Pi^{\star}_{r}(dr)$ a (mixture of) Gaussian processes. The second is to use a scale mixture of Gaussian processes of the form $r\mid\sigma_{r}]\sim\operatorname{GP}(0,\kappa)$ where $\kappa(x,x^{\prime})=\sigma^{2}_{r}\,\rho(x,x^{\prime})$ and $\rho(x,x^{\prime})$ is a fixed kernel and $\sigma^{2}_{r}$ is given a hyperprior. For spike-and-slab priors, Corollary 1 applies immediately.
Applying Theorem 2 directly requires us to have a lower bound on the rate of convergence for Gaussian process models. In other words, we need to verify that a pure GP prior without a spike-and-slab component does not achieve the parametric $N^{-1/2}$ rate. To keep technical details to a minimum, we find it prudent to adopt $\Pi_{\alpha}\{H(\theta_{0},\theta)>M_{N}\,N^{-1/2}\mid\boldsymbol{Z}_{N}\}\to 0$ in $F_{\theta_{0}}$-probability as a starting assumption rather than to state conditions for it to hold. This will almost always be the case when $\kappa(x,x^{\prime})$ is a strictly positive definite kernel such as the squared exponential and Matérn kernels.
Remark 3 (Gaussian Process Lower Bounds).
We might expect flexible nonparametric methods to attain a rate of the form $\epsilon_{N}=N^{-p/(2p+D)}$ for some $p>0$. For some common kernels (such as the squared exponential kernel), however, we only miss $\epsilon_{N}=N^{-1/2}$ by a logarithmic factor (see Theorem 10 of van der Vaart and van Zanten,, 2011). This occurs because the reproducing kernel Hilbert space (RKHS) of the squared exponential kernel consists of extremely smooth functions, making it particularly well suited to approximating the function $r_{0}(x)=0$. The general problem of lower bounding posterior contraction rates for Gaussian process regression models is studied by Castillo, (2008).
If the goal is inference on $\beta_{0}$, it can also suffice to not apply shrinkage and instead project the GP posterior onto the family $\mathcal{F}_{0}$. Recall that using projections for inference is what we actually recommend. For Model NR this is particularly simple due to conjugacy.
Theorem 6 (Semiparametric Bernstein-von Mises Theorem).
Consider Model NR with $\sigma_{0}=1$ and $r_{0}(x)=0$. Assume $\Sigma_{X}=\mathbb{E}_{\theta_{0}}(X_{i}X_{i}^{\top})$ exists, is finite, and is non-singular. Let $\Pi(\cdot\mid\boldsymbol{Z}_{N})$ be the posterior associated to the model with $Y_{i}\mid X_{i},\mu]\sim\operatorname{Normal}\{\mu(X_{i}),1\}$ and $\mu\sim\operatorname{GP}(0,\kappa)$ where $\kappa(x,x^{\prime})=\sigma^{2}_{\beta}\,x^{\top}x^{\prime}+a\,e^{-\rho\|x-x^{\prime}\|^{2}_{2}}$ for some positive constants $(\sigma^{2}_{\beta},a,\rho)$. Let $\beta^{\star}=\arg\min_{\beta}\|\boldsymbol{\mu}-\boldsymbol{X}\beta\|$ where $\boldsymbol{\mu}=(\mu(X_{1}),\ldots,\mu(X_{N}))^{\top}$ and $\boldsymbol{X}=(X_{1},\ldots,X_{N})^{\top}$. Then
$d\left\{\sqrt{N}(\beta^{\star}-\widehat{\beta}_{LS}),\operatorname{Normal}(0,\Sigma_{X}^{-1})\right\}\to 0\quad\text{in $F_{\theta_{0}}$-probability}$
where $d(\cdot,\cdot)$ metrizes convergence in distribution and $\widehat{\beta}_{LS}$ is the least-squares estimator of $\beta_{0}$.
Remark 4.
For a Gaussian process prior with Model NR, the posterior distribution of $\beta^{\star}$ is normal, and so Theorem 6 really only concerns the behavior of the posterior mean and variance. We note that Theorem 6 does not directly rule out that a Bernstein-von Mises result holds without the linear term $\sigma^{2}_{\beta}x^{\top}x^{\prime}$ in the kernel. In Section 4.3 we confirm empirically that Theorem 6 accurately predicts performance, while kernels without the linear term are empirically very highly biased in finite samples.
Theorem 6 is proven using the conjugacy of the GP prior. To show that the result is general, we also prove an analogous result for logistic regression (Model NC). The proof follows Theorem 2.2 of L’Huillier et al., (2023), which is itself based on Theorem 2.1 of Castillo and Rousseau, (2015). The main ingredients of the proof are that (i) $\|r-r_{0}\|=o(N^{-1/4})$ with high posterior probability, which allows us to control the remainder of a Taylor expansion and (ii) $r_{0}(x)=x^{\top}\beta_{0}$ lies in the RKHS of the prior because of the linear term in the kernel.
Theorem 7 (Logistic GP Bernstein–von Mises).
Consider Model NC with $r_{0}(x)\equiv 0$ and $X_{1},X_{2},\ldots$ a deterministic sequence in $0,1]^{P}$. Suppose $\widehat{I}_{N}=N^{-1}\sum_{i}\mu_{0}(X_{i})\{1-\mu_{0}(X_{i})\}X_{i}X_{i}^{\top}$ is invertible and $\widehat{I}_{N}\to I$ for some positive-definite $P\times P$ matrix $I$. Let $\Pi(\cdot\mid\boldsymbol{Z}_{N})$ be the posterior distribution of $r(\cdot)$ associate to the model with $Y_{i}\mid X_{i},r]\sim\operatorname{Bernoulli}\{(1+e^{-r(X_{i})})^{-1}\}$ and $r\sim\operatorname{GP}(0,\kappa)$ where $\kappa(x,x^{\prime})=\sigma_{\beta}^{2}x^{\top}x^{\prime}+\kappa^{\star}(x,x^{\prime})$ for $\sigma_{\beta}^{2}>0$ and $\kappa^{\star}$ satisfying Assumption A in the Supplementary Material. Define $\beta^{\star}=\widehat{I}_{N}^{-1}N^{-1}\sum_{i}\mu_{0}(X_{i})\{1-\mu_{0}(X_{i})\}X_{i}r(X_{i})$. Then
$d\left\{\sqrt{N}(\beta^{\star}-\widehat{\beta}),\operatorname{Normal}(0,I^{-1})\right\}\to 0\quad\text{in $F_{\theta_{0}}$-probability}$
where $d(\cdot,\cdot)$ metrizes convergence in distribution and $\widehat{\beta}=\beta_{0}+N^{-1}\widehat{I}_{N}^{-1}\sum_{i=1}^{N}\{Y_{i}-\mu_{0}(X_{i})\}X_{i}$.
Theorem 7 shows that a certain “information-weighted” projection $\beta^{\star}$ satisfies a semiparametric Bernstein–von Mises theorem with $I$ the efficient Fisher information under $\mathcal{F}_{0}$. The quantity $\widehat{\beta}$ is equivalent to the MLE under the parametric logistic regression submodel, being a one-step Newton–Raphson update of a $\sqrt{N}$-consistent quantity. We use the information-weighted projection mainly for technical convenience, since it makes $\beta^{\star}$ a linear functional of $r(\cdot)$. In practice, it is much better to use the Kullback-Leibler projection $\beta^{\star}=\arg\min_{\beta}\sum_{i}\mu(X_{i})\log\mu(X_{i})/\mu_{\beta}(X_{i})$ where $\mu_{\beta}(x)=(1+e^{-x^{\top}\beta})^{-1}$, as in experiments we have found it to be far less biased.
4 Applications
4.1 Rate Adaptivity of General BART
We now show empirically that the GBART model satisfies our rate adaptivity result. In particular (i) it performs just as well as a correctly specified linear regression model when the linear model holds, and (ii) it performs as well as BART when the linear model is misspecified. We consider the data generating mechanism
$Y_{i}=\frac{1}{\sqrt{P}}\sum_{j=1}^{P}X_{ij}+\lambda_{0}\,X_{i1}^{2}+\epsilon_{i},\quad\epsilon_{i}\sim\operatorname{Normal}(0,\sigma_{0}^{2}),$
with $X_{ij}\sim\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\operatorname{Normal}(0,1)$. When $\lambda_{0}=0$ the model is exactly linear, while $\lambda_{0}\neq 0$ introduces a quadratic component. We consider $\sigma_{0}\in\{1,3,5\}$ and $\lambda_{0}\in\{0,0.4\}$. For each $N\in\{2^{6},2^{7},\ldots,2^{12}\}$ we simulate five datasets and compare predictive accuracy of the following models:
  • BART: the original BART model of Chipman et al., (2010) with its default prior.
  • GBART: the general BART model with its default prior and a flat prior on $\beta$.
  • Linear: a parametric linear regression fit via least squares.
Performance Metrics
Methods are compared by mean squared error (MSE) in estimating $\mu_{0}(x)=\mathbb{E}_{\theta_{0}}(Y_{i}\mid X_{i}=x)$. The MSE is $\frac{1}{N}\sum_{i}\{\mu_{0}(X_{i})-\widehat{\mu}(X_{i})\}^{2}$, where $\widehat{\mu}(x)$ is the posterior mean of $\mu(x)$.
Figure 2: Results for the simulation experiment in Section 4.1. From left to right, we have $\sigma_{0}=1,3,5$ and from top to bottom we have $\lambda_{0}=0,0.4$. MSE and $N$ are displayed on the log-scale.
Results
Figure 2 shows the results. We see that when $\lambda_{0}=0$, the performance of GBART is indistinguishable from that of the linear model at all sample sizes. Somewhat surprisingly, for smaller signal-to-noise ratios and sample sizes the unmodified BART model outperforms a correctly-specified linear model; this is because the sample size is not large enough to estimate the coefficients reliably, giving an advantage to models that regularize the predictions towards zero. This defect in the linear model can be mitigated by replacing the least squares estimates with a ridge regression or lasso.
When $\lambda_{0}=0.4$, GBART outperforms the other methods, particularly at large sample sizes. It is strictly superior to the linear model under all data generating mechanisms. We again see that when both the sample size and the signal-to-noise ratios are small there is an advantage to using a BART model that more strongly regularizes the predictions.
4.2 Adaptivity and Selection of a Gaussian Process
We also confirm empirically that the spike-and-GP prior can consistently select the correct parametric submodel. We use the same simulation design as in Section 4.1. For the prior on $r(x)$ we take $r\sim\operatorname{GP}(0,\sigma_{\mu}^{2}\kappa)$ with $\sigma_{\mu}^{2}\sim p_{0}\,\delta_{0}+(1-p_{0})\operatorname{InvGam}(a_{\sigma_{\mu}},b_{\sigma_{\mu}})$ and $\kappa(x,x^{\prime})=\exp\{-\rho\|x-x^{\prime}\|^{2}\}$ with $\rho\sim\operatorname{InvGam}(a_{\rho},b_{\rho})$. In our illustrations we set $a_{\sigma_{\mu}}=b_{\sigma_{\mu}}=a_{\rho}=b_{\rho}=1$. The inverse-gamma prior was chosen because it puts little mass near zero, so that the model is genuinely nonparametric when $r(x)\neq 0$. We set $p_{0}=1/2$ to express prior indifference between parametric and nonparametric models, though in practice one might prefer a smaller $p_{0}$.
Performance Metrics
We focus on the posterior probability assigned to the nonparametric model, $\Pi(r\neq 0\mid\boldsymbol{Z}_{N})$. When $\lambda_{0}=0$ (linear model), we want $\Pi(r\neq 0\mid\boldsymbol{Z}_{N})\to 0$, while when $\lambda_{0}\neq 0$ we want $\Pi(r\neq 0\mid\boldsymbol{Z}_{N})\to 1$.
Figure 3: Posterior inclusion probabilities for the Gaussian process model under various values of $\lambda_{0}$. From left to right: $\sigma_{0}\in\{1,2,4\}$. From top to bottom: $N\in\{200,400,800\}$.
Results
Figure 3 shows that the spike-and-GP prior behaves as expected. When $\lambda_{0}=0$ the parametric model is consistently selected, while as $\lambda_{0}$ increases the nonparametric model is selected with high probability. The trend is more muted when $\sigma_{0}$ is larger; this is intuitive, since (for example) $\sigma_{0}=4$ corresponds to a signal-to-noise ratio of roughly 16-to-1, while $\sigma_{0}=1$ corresponds to a signal-to-noise ratio of roughly 1-to-1.
As a general lesson, these results suggest that projection-based inference is often preferable to relying solely on model selection consistency. Projections allow us to quantify the loss from using the parametric approximation directly, which can be more informative than a binary model selection outcome. We examine this in more detail in Section 4.3.
4.3 Semiparametric Bernstein–von Mises for Gaussian Processes
We now empirically validate the Bernstein–von Mises result for the projection of the GP posterior onto the linear regression model. We use the data generating process
$Y_{i}=\sum_{j=1}^{P}\beta_{0j}X_{ij}+\epsilon_{i},$
with $X_{ij}\sim\operatorname{Normal}(0,1)$ and $\epsilon_{i}\sim\operatorname{Normal}(0,1)$ independently. We set $P=5$ and $\beta_{0j}\equiv 1.55$. We then fit a GP regression model $Y_{i}=\mu(X_{i})+\epsilon_{i}$ with $\mu\sim\operatorname{GP}(0,\kappa)$ and compute the posterior distribution of the projection parameter $\beta^{\star}=(\boldsymbol{X}^{\top}\boldsymbol{X})^{-1}\boldsymbol{X}^{\top}\boldsymbol{\mu}$. We consider the Laplace kernel $\kappa(x,x^{\prime})=e^{-\|x-x^{\prime}\|}$, the squared exponential kernel $\kappa(x,x^{\prime})=e^{-\|x-x^{\prime}\|^{2}}$, and the squared-exponential-plus-linear kernel $\kappa(x,x^{\prime})=100\,(x^{\prime})^{\top}x+e^{-\|x-x^{\prime}\|_{2}}$.
Figure 4: Results for the experiments in Section 4.3.
Performance Metrics
We monitor three quantities: (i) the nominal coverage of 95% credible intervals for $\beta^{\star}_{1}$ under each kernel; (ii) the scaled bias of the posterior mean, defined as the average of $\sqrt{N}(\widehat{\beta}_{1}-\beta_{01})$ with $\widehat{\beta}=\mathbb{E}_{\Pi}(\beta^{\star}\mid\boldsymbol{Z}_{N})$; and (iii) the scaled variance, defined as the average of $N\operatorname{Var}_{\Pi}(\beta^{\star}_{1}\mid\boldsymbol{Z}_{N})$. If the posterior satisfies a semiparametric Bernstein–von Mises result, we expect coverage near 95%, scaled bias near zero, and scaled variance near 1, since $\mathbb{E}_{\theta_{0}}(\boldsymbol{X}^{\top}\boldsymbol{X})=\mathrm{I}$ in this setup.
Results
Figure 4 shows results across kernels and sample sizes. Among the three, only the squared-exponential-plus-linear kernel empirically satisfies the semiparametric Bernstein–von Mises theorem at the chosen sample sizes. In larger samples, we do see the asymptotic variance of the squared exponential and Laplace kernels approaching the semiparametric efficiency, and indeed asymptotically we do attain the semiparametric efficiency level; the problem with these kernels is that, even in very large samples, the estimates are highly biased, which is clearly seen in the middle panel of Figure 4.
4.4 Diagnosing Model Misspecification in Practice
We now show how to diagnose model misspecification in practice by analyzing data from the Medical Expenditure Panel Survey (MEPS, Cohen et al.,, 2009). We consider a subset of the data from the 2011 survey containing 16,113 with the goal of predicting self-assessed health status. We use Model NR with $Y_{i}$ representing the self-assessed health status of individual $i$ and $X_{i}$ representing the individual’s age, body mass index (BMI), education level, income relative to the poverty line, region of the country, sex, marital status, race, seatbelt use, and smoking status.
To illustrate the ability of the GBART model to default to a linear model when there is little data while also adapting to non-linearities and interactions in the data, we consider two analyses: one using a randomly selected sample of size of $N=1000$ and one using a sample of size $N=8000$. For both cases we monitor the summary $R^{2}$ of the projection onto a linear regression. To assess how closely the semiparametric Bernstein-von Mises result holds, we also monitor the posterior distribution of the coefficient of BMI.
Figure 5: For $N=1000$: the posterior summary $R^{2}$ and overall summary $R^{2}$ of the linear model (left), posterior distribution of the coefficient of BMI compared with the result from the linear model (middle), and plot of the projected predictions against the actual predictions from the model (right).
Results for the $N=1000$ model are given in Figure 5. We see that the model is shrunk very heavily towards a linear regression model, with the posterior summary $R^{2}$ being very close to $100\%$. The posterior distribution of the coefficient of BMI is also very close to the posterior from a parametric Bayesian linear model, indicating that the semiparametric Bernstein-von Mises approximation is accurate. Finally, the projected linear regression model gives predictions that are virtually indistinguishable from the full general BART model. This indicates that, at $N=1000$, there is limited evidence of model misspecification so that the general BART model defaults to the linear model.
Figure 6: For $N=8000$: the posterior summary $R^{2}$ and overall summary $R^{2}$ of the linear model (left), posterior distribution of the coefficient of BMI compared with the result from the linear model (middle), and plot of the projected predictions against the actual predictions from the model (right).
By contrast, results for the $N=8000$ model are given in Figure 6. While the summary $R^{2}$ of the model is still quite high, we see that the linear regression model no longer captures all of the variability in the predictions. Nevertheless, the semiparametric Bernstein-von Mises result still appears to hold, with the projection parameter $\beta_{\text{BMI}}$ having a posterior distribution in agreement with the parametric linear regression.
Figure 7: For the $N=8000$ additive model: the posterior summary $R^{2}$ and overall summary $R^{2}$ of the linear model (left) and plot of the projected predictions against the actual predictions from the model (right).
From the analysis for the $N=8000$ setting, we see that there is evidence of misspecification of the linear regression, implying that there exist either non-linearities or interactions in the data. An advantage of the fully-Bayesian approach is that we are free to summarize the posterior in any way we choose, and in particular we can look for better models to project the posterior of $\mu(x)$ on to. To this end, in Figure 7 we give the summary $R^{2}$ and projected predictions for a model that includes spline terms for age, BMI, poverty level, with interaction terms for (age) $\times$ BMI and (age) $\times$ (poverty level). After making this modification, we find that the overall summary $R^{2}$ of the model exceeds 98%, with the posterior distribution of $R^{2}$ centered around 96%, giving a roughly $10\%$ increase in predictive variability accounted for.
Figure 8: Decision tree fit to the residuals $\widehat{\mu}(X_{0})-\widetilde{\mu}(X_{i})$ from the additive model.
To gain better insight into the remaining sources of variability, Figure 8 displays a CART fit to the residuals $\widehat{\mu}(X_{i})-\widetilde{\mu}(X_{i})$, where $\widetilde{\mu}(x)$ denotes the additive summary of $\widehat{\mu}(x)$. The magnitude of the predictions of the CART model represent how off, on average, the additive summary is from the general BART model for subsets of the observations. The largest degree of misfit occurs for individuals who are white and with poverty level less than $121$, with the additive summary predicting a health status on average 0.036 units higher than the GBART model.
Summary
Overall, we see that the general BART model is able to adapt to the amount of evidence in favor of the linear model, reverting to the linear model when there is little evidence of misspecification. In addition, when there is evidence of misspecification, we have tools for inspecting the posterior to identify subsets of the data that are responsible for the misspecification. We feel that the additive summary is adequate (0.036 off in a small subgroup, relative to a residual standard deviation of around 1), although we could go further in this example by adding, for example, an interaction between poverty level and race.
5 Discussion
The position advocated for in this manuscript is one of defensive model expansion, wherein we embed a parametric Bayesian model inside a nonparametric Bayesian model and apply strong shrinkage towards the parametric model. We claim that this retains many of the benefits of specifying a parametric Bayesian model, such as strong performance when $N$ is small or the signal-to-noise ratio is low, while mitigating the drawbacks of potential model misspecification.
The principal drawback of defensive model expansion is that the expanded nonparametric model may be less interpretable than the embedded parametric model. For example, how should we interpret $\eta$ if the nonparametric component is active in 50% of posterior samples but inactive in the other 50%? On this point, an advantage of defensive model expansion is that we also obtain measures of the degree of misspecification. If the summary $R^{2}$ of the parametric submodel is sufficiently large, we may reasonably conclude that little is lost by interpreting the parametric submodel directly, together with guidelines about “how wrong” it is expected to be. We find the Bayesian nonparametric approach particularly appealing for implementing this strategy because it can be carried out coherently through posterior projections, without requiring the analyst to fit alternative models. In this way, defensive model expansion provides both robustness to misspecification and a principled path for interpretation.

### v2

@@
 Some applied researchers hesitate to use nonparametric methods, worrying that they will lose power in small samples or overfit the data when simpler models are sufficient.
-We argue that at least some of these concerns are unfounded when nonparametric models are strongly shrunk towards parametric submodels.
+We argue that at least some of these concerns are unfounded when nonparametric models are strongly shrunk toward parametric submodels.
 We consider expanding a parametric model with a nonparametric component $r(x)$ that is heavily shrunk toward zero.
@@
 We show that this adaptive behavior follows from simple and general conditions.
-Specifically, we prove that Bayesian nonparametric models anchored to linear regression, including variants of Gaussian processes regression and Bayesian additive regression trees, consistently identify the correct parametric submodel when it holds and give asymptotically efficient inference for regression coefficients.
-In simulations, we find that the general BART model performs identically to correctly specified linear regression when the parametric model holds, and substantially outperform it when nonlinear effects are present.
+Specifically, we prove that Bayesian nonparametric models anchored to linear regression, including variants of Gaussian process regression and Bayesian additive regression trees, consistently identify the correct parametric submodel when it holds and give asymptotically efficient inference for regression coefficients.
+In simulations, we find that the general BART model performs identically to correctly specified linear regression when the parametric model holds, and substantially outperforms it when nonlinear effects are present.
 This suggests a practical paradigm: defensive model expansion as a safeguard against model misspecification.
@@
   • If the parametric model is badly misspecified or omits important signals, expanding into a nonparametric model can improve power.
-With a flexible “alternative” model where $r(x)\neq 0$, we can also use the estimated size of $r(\cdot)$ as a diagnostic for parametric model misspecification: if $|r(x)|$ is large on average then this is evidence that the parametric submodel model fits poorly.
+With a flexible “alternative” model where $r(x)\neq 0$, we can also use the estimated size of $r(\cdot)$ as a diagnostic for parametric model misspecification: if $|r(x)|$ is large on average then this is evidence that the parametric submodel fits poorly.
   • If the nonparametric component turns out to be important, we obtain more honest uncertainty quantification and the chance for richer scientific discoveries.
@@
 This idea is a core theme in the Bayesian nonparametric literature, although in practice much of the focus has been on clustering and latent feature models (Orbanz and Teh,, 2011; Gershman and Blei,, 2012; Ghahramani et al.,, 2007).
-A key difference from the settings we consider is that clustering and latent feature problems are evidently complex (with the main question being only how complex), whereas we focus on settings that are plausibly simple, where a parametric model might in fact be sufficient.
+A difference from the settings we consider is that clustering and latent feature problems are evidently complex (with the main question being only how complex), whereas we focus on settings that are plausibly simple, where a parametric model might in fact be sufficient.
 See Li and Dunson, (2020) for an example of using a baseline nonparametric model as a tool for evaluating parametric models, as we do here.
@@
 Neal, (1995) makes the point that, “There is no provision in the Bayesian framework for changing the model or the prior depending on how much data was collected.
-If the model and prior are correct for a thousand observations, they are correct for ten observations as well (though the impact of using an incorrect prior might be more serious with fewer observations),” and that when the underlying phenomena is evidently complex the appropriate solution to finding a simpler model outperforming a more complex model is “to design a different complex model that captures whatever aspect of the problem led to the simple model performing well.”
+If the model and prior are correct for a thousand observations, they are correct for ten observations as well (though the impact of using an incorrect prior might be more serious with fewer observations),” and that when the underlying phenomenon is evidently complex the appropriate solution to finding a simpler model outperforming a more complex model is “to design a different complex model that captures whatever aspect of the problem led to the simple model performing well.”
 In the context of mixture models, a selling point of infinite mixtures $f(y)=\sum_{k=1}^{\infty}\omega_{k}\,g(y\mid\vartheta_{k})$ is that they can collapse to a single parametric model $g(y\mid\theta)$ when most prior mass is placed on the event $\omega_{1}\approx 1$.
-For Dirichlet process mixtures (Escobar and West,, 1995), this is accomplished by shrinking the concentration parameter heavily towards zero.
+For Dirichlet process mixtures (Escobar and West,, 1995), this is accomplished by shrinking the concentration parameter heavily toward zero.
 For conditional mixture models — such as dependent Dirichlet processes (DDPs, MacEachern,, 2000; Quintana et al.,, 2022) — this allows us to recover standard parametric generalized linear models when appropriate (Shahbaba and Neal,, 2009; Hannah et al.,, 2011).
 Our strategy also represents a type of continuous model expansion in the sense of Gelman and Shalizi, (2013), who recommend “forming a larger model that includes both A and B as special cases,” as an alternative to testing Model A against Model B.
-Like them, we do not formally test the parametric submodel (except as a theoretical device).
+Like them, we do not recommend formally testing the parametric submodel (except as a theoretical device).
 While they explicitly do not recommend Bayesian nonparametrics as a tool for continuous model expansion, we also do not feel that our proposal is at odds with the additional use of model checking they recommend, and do not view our use of an additional nonparametric component to be a complete solution to model misspecification.
 Related works include Kamary et al., (2018) and Yao et al., (2018).
-1.2 Contributions
+1.2 Outline
 We recommend defensive model expansion partly because, under very general conditions, it yields efficient inference when the associated parametric model is correctly specified.
@@
 In fact, the nonparametric component of the prior plays little direct role in our analysis, making the results extremely general.
+That being said, the main role played by the fractional posterior is to make establishing a posterior contraction rate simpler; the assumption that a fractional posterior is used can often be replaced with directly assuming a posterior concentration rate, for which there are many existing rules for the GP and BART models we consider in this work.
 We begin by presenting general results on posterior concentration for nonparametric models when parametric submodels have sufficient prior mass.
-With the exception of Theorems 6 and 7, these are straightforward extensions of existing results for fractional posteriors.
+These are straightforward extensions of existing results for fractional posteriors.
 We prove results for convergence rates, model selection, and asymptotic normality under the assumption that the parametric model is correctly specified.
-Our model selection results essentially show that (i) we can correctly select the parametric model and (ii) the posterior distribution satisfies a Bernstein–von Mises theorem.
-We also prove two semiparametric Bernstein–von Mises theorems that apply in settings where model selection consistency is difficult to use, which guarantees efficient posterior concentration of a projection parameter $\theta^{\star}$ around its true value.
+Our model selection results essentially show that (i) we can correctly select the parametric model and (ii) the posterior distribution satisfies a Bernstein-von Mises theorem.
+We also prove two semiparametric Bernstein-von Mises theorems (Theorem 6 and Theorem 7) that apply in settings where model selection consistency is difficult to apply, which guarantee efficient posterior concentration of a projection parameter $\theta^{\star}$ around its true value; when the submodel is correct, this projection parameter will coincide with the parameter of interest.
 Such projection parameters are often used to summarize posterior distributions (Woody et al.,, 2021).
-We then apply these results to Gaussian process (GP) and Bayesian additive regression tree (BART) models.
-We show that BART models naturally satisfy the conditions required for a Bernstein–von Mises result (see also Rocková,, 2020), as do “spike-and-GP” priors.
-While we do not study this here, our semiparametric Bernstein–von Mises results for GPs also have scope for being extended to projection parameters even when the anchoring parametric model is misspecified.
+To the best of our knowledge, Theorem 7 is not implied by existing results.
+We apply these results to Gaussian process (GP) and Bayesian additive regression tree (BART) models.
+We show that BART models naturally satisfy the conditions required for a Bernstein-von Mises result (see also Rocková,, 2020), as do “spike-and-GP” priors.
+While we do not study this here, our semiparametric Bernstein-von Mises results also may be extended to projection parameters even when the anchoring parametric model is misspecified.
 These theoretical results are validated on simulated data.
 In addition to being theoretically simple, defensive model expansion is easy to implement in practice.
-For example, because our convergence rate and model selection results apply directly to BART models, existing software for the general BART model of Tan and Roy, (2019) makes it straight-forward to obtain parametric adaptivity with default prior specifications.
+For example, because our convergence rate and model selection results apply directly to BART models, existing software for the general BART model of Tan and Roy, (2019) makes it straightforward to obtain parametric adaptivity with default prior specifications.
 BART models are especially attractive because, in addition to adapting to parametric structures, they also adjust to locally varying smoothness and the presence of low-order interactions (Linero and Yang,, 2018; Ročková and van der Pas,, 2020; Jeong and Rockova,, 2023).
@@
 For any function $g$, let $\mathbb{E}_{\theta}\{g(Z)\}$ and $\operatorname{Var}_{\theta}\{g(Z)\}$ denote expectation and variance under $f_{\theta}$.
-We assume that $\theta$ decomposes as $(r,\eta)$ where $r\in\mathcal{H}$ is infinite-dimensional and $\eta\in E\subseteq\mathbb{R}^{P}$ is-finite dimensional, such that $\Theta=\mathcal{H}\times E$.
+We assume that $\theta$ decomposes as $(r,\eta)$ where $r\in\mathcal{H}$ is infinite-dimensional and $\eta\in E\subseteq\mathbb{R}^{P}$ is finite-dimensional, such that $\Theta=\mathcal{H}\times E$.
 The family $\mathcal{F}_{0}=\{F_{\theta}:\eta\in E\ \text{and}\ r=r_{0}\}$ represents a parametric submodel of $\mathcal{F}$.
-In our examples, $r(\cdot)$ is a mapping $\mathbb{R}^{P}\to\mathbb{R}$ and $r_{0}(\cdot)\equiv 0$.
+In our examples, $r(\cdot)$ is a mapping $\mathbb{R}^{D}\to\mathbb{R}$ and $r_{0}(\cdot)\equiv 0$.
 Intuitively, $\eta$ corresponds to the usual parametric component (e.g., regression coefficients), while $r$ is a flexible nonparametric adjustment.
@@
 Here $K(\theta_{0}\|\theta)$ and $V(\theta_{0}\|\theta)$ are Kullback–Leibler type divergences, as defined in Section 3.
+Different examples where we apply PPA will use different norms and inner products.
 We now describe three concrete models where our results apply. We will focus mainly on Model NR for clarity.
@@
 Anchoring to the parametric submodel with $r_{0}(x)\equiv 0$ recovers ordinary linear regression.
-Our goals are to estimate the regression function $\mu_{0}(x)=x^{\top}\beta_{0}$ and, ideally, recover $\beta_{0}$ efficiently.
+Our goals are to estimate the regression function $\mu_{0}(x)=r_{0}(x)+x^{\top}\beta_{0}$ and, ideally, recover $\beta_{0}$ efficiently.
 Nonparametric Classification (Model NC)
-We adopt the same framework as Model NR, except $Y_{i}\in\{0,1\}$ and $\mathbb{E}_{\theta}(Y_{i}\mid X_{i}=x)=\mu(x)=(1+\exp\{x^{\top}\beta+r(x)\})^{-1}$ so that $Y_{i}\mid X_{i}=x]$ is a logistic regression and $\theta=(r,\beta)$.
-Our goal again is to estimate $\mu_{0}(x)$ and estimate $\beta_{0}$ efficiently.
+We adopt the same framework as Model NR, except $Y_{i}\in\{0,1\}$ and $\mathbb{E}_{\theta}(Y_{i}\mid X_{i}=x)=\mu(x)=(1+\exp\{-x^{\top}\beta-r(x)\})^{-1}$ so that $Y_{i}\mid X_{i}=x]$ is a logistic regression and $\theta=(r,\beta)$.
+Our goal again is to estimate $\mu_{0}(x)=\mathbb{E}_{\theta_{0}}(Y_{i}\mid X_{i}=x)$ and estimate $\beta_{0}$ efficiently.
 Density Regression (Model DR)
@@
 For routine use, we recommend the Bayesian Additive Regression Trees (BART, Chipman et al.,, 2010) prior for the nonparametric component $r(\cdot)$.
-In the context of Model NR the regression function is $\mu(x)=x^{\top}\beta+\sum_{t=1}^{T}g(x;\mathcal{T}_{t},\mathcal{M}_{t})$ where each $g(x;\mathcal{T}_{t},\mathcal{M}_{t})$ is a regression tree, $\mathcal{T}_{t}$ denotes the tree structure, and $\mathcal{M}_{t}=\{\lambda_{\ell t}:\ell=1,\ldots,L_{t}\}$ is the set of leaf node predictions.
+In the context of Model NR, the regression function is $\mu(x)=x^{\top}\beta+\sum_{t=1}^{T}g(x;\mathcal{T}_{t},\mathcal{M}_{t})$ where each $g(x;\mathcal{T}_{t},\mathcal{M}_{t})$ is a regression tree, $\mathcal{T}_{t}$ denotes the tree structure, and $\mathcal{M}_{t}=\{\lambda_{\ell t}:\ell=1,\ldots,L_{t}\}$ is the set of leaf node predictions.
 Tan and Roy, (2019) refer to this as a general BART (GBART) model, and it can be fit using the SoftBart package (Linero,, 2022) using the gsoftbart_regression function.
 A schematic of the construction of the model is given in Figure 1, which shows a single tree from the ensemble and the associated step function.
-Figure 1: Left: an example of a regression tree with input $x=(x_{1},x_{2})$ supported on $[0,1]$. Right: the induced step function on $0,1]^{2}$ .
+Figure 1: Left: an example of a regression tree with input $x=(x_{1},x_{2})$ supported on $0,1]^{2}$. Right: the induced step function on $0,1]^{2}$ .
 We now review the prior specification of BART, which is important for our purposes.
 The default recommendation of Chipman et al., (2010) takes $\lambda_{t\ell}\sim\operatorname{Normal}(0,\sigma^{2}_{\mu}/T)$ with $\sigma_{\mu}=0.5$ after centering and scaling the $Y_{i}$’s to lie between -0.5 and 0.5.
-The scaling of $\sigma^{2}_{\mu}$ by $T$ is done so that $\operatorname{Var}\{\sum_{t}g(x;\mathcal{T}_{t},\mathcal{M}_{t})\}=\sigma^{2}_{\mu}$, which in turn leads to a Gaussian process limit as $T\to\infty$ (Linero,, 2017).
+The scaling of $\sigma^{2}_{\mu}$ by $T$ is done so that $\operatorname{Var}\{\sum_{t=1}^{T}g(x;\mathcal{T}_{t},\mathcal{M}_{t})\}=\sigma^{2}_{\mu}$, which in turn leads to a Gaussian process limit as $T\to\infty$ (Linero,, 2017).
 One way to shrink the GBART model to a linear regression is to place a prior on $\sigma_{\mu}$ supported near 0, such as a half-normal or half-Cauchy prior.
@@
 In Section 3.2 we will show that this suffices for GBART to possess strong theoretical properties when the linear model is correct.
-Chipman et al., (2010) propose a branching process prior for $\mathcal{T}_{t}$ that (i) decides whether each node at a given depth of the tree will be converted into a branch node with two leaf children; (ii) iterating over the depths; and (iii) terminating if all of the nodes at a given depth are leaf nodes.
+Chipman et al., (2010) propose a branching process prior for $\mathcal{T}_{t}$ that (i) decides whether each node at a given depth of the tree will be converted into a branch node with two leaf children; (ii) iterates over the depths; and (iii) terminates if all of the nodes at a given depth are leaf nodes.
 The probability of a node at depth $d$ being a branch node under the prior is specified to be $q(d)=a/(1+b)^{d}$ where by default $a=0.95$ and $b=2$.
-In particular, the probability that a tree consists of no splits is (1a)(1-a), and the prior probability that the GBART model reduces to the linear model is therefore $(1-a)^{T}>0$.
+In particular, the probability that a tree consists of no splits is (1a)(1-a), and the prior probability that the GBART model reduces to the linear model (with intercept offset by the constant $r(0)$) is therefore $(1-a)^{T}>0$.
 2.2 Gaussian Process Regression
 A common approach for anchoring Model NR to a parametric submodel is to use the Gaussian process regression (GP regression) model.
-GPs are distributions on random functions, and can be used as priors model for regression functions $r(\cdot)$.
+GPs are distributions on random functions, and can be used as priors for regression functions $r(\cdot)$.
 A GP is determined by a mean function $m(x)$ and a positive semi-definite kernel $\kappa(x,x^{\prime})$, with $\mathbb{E}_{\Pi}r(x)=m(x)$ and $\operatorname{Cov}_{\Pi}\{r(x),r(x^{\prime})\}=\kappa(x,x^{\prime})$.
@@
 Equivalently, letting $\mu(x)=x^{\top}\beta+r(x)$ we have $\mu\sim\operatorname{GP}(m,\sigma^{2}_{\mu}\,\kappa)$ where $m(x)=x^{\top}\beta$.
-If $\beta\sim\operatorname{Normal}(0,\sigma^{2}_{\beta})$ this simplifies further to $\mu\sim\operatorname{GP}(0,\sigma^{2}_{\mu}\kappa+\sigma^{2}_{\beta}\kappa_{\text{linear}})$ where $\kappa_{\text{linear}}(x,x^{\prime})=x^{\top}x^{\prime}$.
+If $\beta\sim\operatorname{Normal}(0,\sigma^{2}_{\beta}\mathrm{I})$ this simplifies further to $\mu\sim\operatorname{GP}(0,\sigma^{2}_{\mu}\kappa+\sigma^{2}_{\beta}\kappa_{\text{linear}})$ where $\kappa_{\text{linear}}(x,x^{\prime})=x^{\top}x^{\prime}$.
 In our illustrations we use the squared exponential function $\kappa_{\text{SE}}(x,x^{\prime})=\exp(-\rho\|x-x^{\prime}\|^{2})$, with the lengthscale $\rho$ controlling the smoothness of samples from the prior.
@@
 Scale Mixture Models
-One option is to place a hyperprior on $\sigma_{\mu}$, with heavy shrinkage towards zero.
+One option is to place a hyperprior on $\sigma_{\mu}$, with heavy shrinkage toward zero.
 We call this the scale mixture of GPs approach.
@@
 Semiparametric Bernstein-von Mises for Projections
-A third strategy is to simply use a GP for $r(x)$ and then perform inference on a projection parameter $\eta^{\star}$ obtained from projecting the full model onto the parametric submodel $\mathcal{F}_{0}$.
-Concretely, we might compute the Kullback-Leibler projection $\theta^{\star}=\arg\min_{\theta\in\mathcal{F}_{0}}K(\theta\|\theta^{\star})$ for each sample $\theta$ from the posterior and, because $\theta_{0}\in\mathcal{F}_{0}$, hopefully $\theta^{\star}$ will concentrate faster around $\theta_{0}$ than $\theta$.
-This can be used with either of the previous two approaches, but sometimes works on its own as well.
-When the parametric model is correctly specified, appropriately chosen GP priors can yield efficient inference.
-However, poorly chosen priors may still lead to inefficient or invalid inference.
-For example, we will see that even if the posterior converges at a near-parametric rate such as $N^{-1/2}(\log N)^{t}$ for some $t>0$, the bias in estimating $\beta_{0}$ may decay very slowly, making inference inefficient in finite samples.
-This occurs, for example, when we incorrectly set $\beta=0$ and model $r(x)$ with a squared exponential kernel.
+[…]
 2.3 Orthogonalization
@@
 That is, we require that $\sum_{i}r(X_{i})\,X_{i}=\mathbf{0}$.
+Here this is meant primarily as an empirical, finite-sample orthogonalization device based on the observed design matrix rather than as a population-level identifiability assumption.
 This condition serves two purposes.
 First, it can substantially improve the mixing of MCMC algorithms used to fit the model.
-Second, it prevents $r(\cdot)$ from becoming collinear with $X_{i}^{\top}\beta$, making it easier to shrink $r(\cdot)$ toward zero when it is unnecessary.
+Second, it prevents $r(\cdot)$ from becoming colinear with $X_{i}^{\top}\beta$, making it easier to shrink $r(\cdot)$ toward zero when it is unnecessary.
 Orthogonalization is not required for our theoretical results to hold, but it is often useful in practice.
@@
     If the parametric model is inadequate with high posterior probability, fit additional summaries (such as additive models or tree-based fits) to identify specific aspects of the misspecification, and use these to refine $\mathcal{F}_{0}$.
-One practical implementation is to form a linear projection summary by minimizing
-$\text{SSE}=\sum_{i=1}^{N}\{\widehat{\mu}(X_{i})-X_{i}^{\top}\beta^{\star}\}^{2},$ (1)
-where $\widehat{\mu}(x)$ is the posterior mean function (or a posterior draw).
-We then compute $R^{2}=1-\frac{\text{SSE}}{\sum_{i}\{\widehat{\mu}(X_{i})-\bar{\mu}\}^{2}},\text{ where }\bar{\mu}=N^{-1}\sum_{i}\mu(X_{i}).$ If $R^{2}$ is sufficiently close to 1, we conclude that the parametric model is adequate.
+One practical implementation is to form a linear projection summary by minimizing $\text{SSE}=\sum_{i=1}^{N}\{\mu(X_{i})-X_{i}^{\top}\beta^{\star}\}^{2}.$ We then compute $R^{2}=1-\frac{\text{SSE}}{\sum_{i}\{\mu(X_{i})-\bar{\mu}\}^{2}},\text{ where }\bar{\mu}=N^{-1}\sum_{i}\mu(X_{i}).$ We can interpret $R^{2}$ as the proportion of the overall signal that is explained by the parametric model.
+If $R^{2}$ is sufficiently close to 1, we conclude that the parametric model is adequate.
 The required closeness depends on context: if prediction is the main goal, we might demand $R^{2}$ very close to 1, while if the goal is only a broad scientific conclusion, a more modest $R^{2}$ could suffice.
+As a reporting convention, we recommend presenting the posterior mean and posterior distribution of $R^{2}$; for many descriptive purposes, values above roughly $0.99$ indicate an essentially interchangeable summary, values around $0.95$ indicate a close but imperfect summary, and materially smaller values suggest that richer structure may matter.
 If the linear projection is inadequate, we expand to an additive model $\mu(x)\approx\sum_{j}\gamma_{j}(x_{j})$ and, if that is still insufficient, apply a CART fit to residuals to detect missing interaction terms.
@@
 First, we may lack power to detect $r=r_{0}$ even if the model technically allows for it.
-Second, even when $r\neq r_{0}$ with high posterior probability, the parametric submodel may still be practically indistinguishable from the nonparametric model in terms of its predictions.
+Second, even when $r\neq r_{0}$ with high posterior probability, the parametric submodel may still be practically indistinguishable from the nonparametric model in terms of its predictions; this can occur in the GBART model, for example, when a shrinkage prior is placed on $\sigma_{\mu}$.
 A criticism of the above strategy is that it is not specific enough, requiring subjective judgments about whether a summary is “adequate” or not.
 On the contrary, we feel that this is a strength, as it forces the analyst to engage with the question of what purpose the fitted model is supposed to serve and whether it is sufficient for that purpose.
-Judgments, such as what an appropriate tradeoff between interpretability and predictive accuracy are, should chiefly be determined in collaboration with subject matter experts.
+Judgments, such as what an appropriate tradeoff between interpretability and predictive accuracy are, should chiefly be determined in collaboration with stakeholders.
 3 Theoretical Results
 We now establish some basic theoretical properties for the BART and GP models.
-To formalize the idea that a properly anchored Bayesian nonparametric model can perform “just as well” as a parametric submodel we will use arguments based on the rate of convergence of a posterior distribution; the rate will either be used as a justification on its own, or as part of the assumptions for stronger justifications.
-We will say that the posterior convergence rate is (upper bounded by) $\epsilon_{N}$ if we have $\Pi_{\alpha}\{H(\theta_{0},\theta)>M_{N}\,\epsilon_{N}\mid\boldsymbol{Z}_{N}\}\to 0$ in probability for all sequences $M_{N}\uparrow\infty$ (where $M_{N}$ is allowed to diverge arbitrarily slowly).
+The results in this section split into two parts: general fractional-posterior concentration and model-selection results under correct parametric specification, and ordinary-posterior semiparametric Bernstein-von Mises results for projection parameters.
+To formalize the idea that a properly anchored Bayesian nonparametric model can perform “just as well” as a parametric submodel, we will use arguments based on the rate of convergence of a posterior distribution; the rate can be used as a justification on its own, or as part of the assumptions for stronger justifications such as model selection consistency or a Bernstein-von Mises theorem.
+We will say that the posterior convergence rate is (upper bounded by) $\epsilon_{N}$ if we have $\Pi_{\alpha}\{H(\theta_{0},\theta)>M\,\epsilon_{N}\mid\boldsymbol{Z}_{N}\}\to 0$ in probability for some $M>0$.
 Here, $H(\theta_{0},\theta)=[\int\{\sqrt{f_{\theta}(z)}-\sqrt{f_{\theta_{0}}(z)}\}^{2}\ dz]^{1/2}$ denotes the Hellinger distance between $f_{\theta}$ and $f_{\theta_{0}}$.
-For example, we will show that both GBART and the spike-and-GP approaches give $\epsilon_{N}=N^{-1/2}$ rates and that posterior inferences for $\beta_{0}$ are efficient, provided that the parametric submodel is correct.
+For example, we will show that appropriate GBART and spike-and-GP priors give $\epsilon_{N}=N^{-1/2}$ rates (possibly up-to a logarithmic term) and that posterior inferences for $\beta_{0}$ are efficient, provided that the parametric submodel is correct.
 The most important condition governing whether parametric adaptivity is possible is whether or not the prior assigns enough mass to neighborhoods of $\theta_{0}$.
@@
 $\Pi_{\alpha}(d\theta\mid\boldsymbol{Z}_{N})=\frac{L_{N}(\theta)^{\alpha}\ \Pi(d\theta)}{\int L_{N}(\theta)^{\alpha}\ \Pi(d\theta)},\qquad\alpha\in(0,1),$ (1)
-where $L_{N}(\theta)=\prod_{i=1}^{N}f_{\theta}(Z_{i})$ is the likelihood.
+where $L_{N}(\theta)=\prod_{i=1}^{N}f_{\theta}(Z_{i})$ is the likelihood; extending to $\alpha=1$ (i.e., the genuine posterior) is possible for the models we consider here under additional conditions on the prior.
 Works that have studied the properties of fractional posteriors include O’Hagan, (1995); Walker and Hjort, (2001); Bhattacharya et al., (2019); Martin and Walker, (2019).
-Remarkably, under essentially no further assumptions, the prior thickness condition is sufficient to guarantee posterior concentration at the rate $\epsilon_{N}$.
+Under essentially no further assumptions, the prior thickness condition is sufficient to guarantee posterior concentration at the rate $\epsilon_{N}$.
 Section 2.2 of Castillo, (2024) is particularly instructive for understanding why these conditions are sufficient and how the arguments can be extended to non-iid settings.
@@
 This construction downweights parameter values that fit the observed data too well, again limiting overfitting.
-This robustness to overfitting is what allows us to avoid conditions on the size of the effective support required of the full posterior.
-Identifiability
-Prior thickness ensures convergence in Hellinger distance, but does not automatically imply efficient estimation of $\eta_{0}$, since $r(\cdot)$ may be confounded with $\eta$.
-For example, in Models NR and NC, $\mu(\cdot)\to\mu_{0}(\cdot)$ does not necessarily imply $\beta\to\beta_{0}$.
-One way to ensure recovery of $\eta_{0}$ is through a metric identifiability condition such as $\|\eta-\eta_{0}\|\lesssim H(\theta,\theta_{0}),$ which forces $H(\theta,\theta_{0})\to 0$ only if $\eta\to\eta_{0}$ at the same rate.
-In Model NR this can be obtained by imposing $\mathbb{E}_{\theta}\{X_{i}r(X_{i})\}=\mathbf{0}$ and assuming $\mathbb{E}(X_{i}\,X_{i}^{\top})=\Sigma_{X}$ is full-rank.
-Alternatively, consistent selection of the event $r=r_{0}]$ also guarantees efficient inference for $\eta_{0}$.
+This robustness to overfitting is what allows us to avoid conditions on the size of the effective support used, for example, by Ghosal et al., (2000) to study the usual posterior.
 Non-iid Data and Finite Sample Bounds
 It is possible to extend our results to allow for non-iid data by instead defining $B_{\epsilon}$ in terms of $\mathbb{E}_{\theta_{0}}\left(\log\frac{f(Z_{1},\ldots,Z_{N}\mid\theta_{0})}{f(Z_{1},\ldots,Z_{N}\mid\theta)}\right)^{k}\leq N\,\epsilon_{N}^{2}$ for $k\in\{1,2\}$, and similarly redefining the Hellinger distance.
-It is also possible to convert the convergence rate arguments into non-asymptotic arguments.
-Finite sample results are given, for example, by Bhattacharya et al., (2019).
+Another advantage of fractional posteriors is that it is also possible to convert convergence rate arguments into non-asymptotic arguments and allow for model misspecification, as done by Bhattacharya et al., (2019).
 Non-Fractional Posteriors
@@
 In this case, prior thickness alone is not enough; one must also control the complexity of the support of the prior, often by placing high mass on a low-entropy sieve (Ghosal et al.,, 2000).
-Prior support conditions serve to stop overfitting to the data.
-While this strategy works, it places much stronger restrictions on the choice of $\Pi_{r}(dr)$ and can introduce an extra $(\log N)^{t}$ factor in the rates that can be cumbersome to remove.
+While this strategy works, it places much stronger restrictions on the choice of $\Pi_{r}(dr)$.
+Nevertheless, we note that the specific semiparametric Bernstein-von Mises results given in Theorem 5, Theorem 6, and Theorem 7 happen to be true for genuine posteriors; the fractional posterior only makes establishing the required rate of convergence easier, but convergence rate results for the non-fractional posterior are widely available for BART and GP models.
 3.1 General Results
@@
 Our first result is a general fractional posterior concentration theorem, implied for example by Theorem 2.1 of Castillo, (2024).
+It establishes a near-parametric rate of convergence under essentially only a prior thickness condition.
 Theorem 1 (General Rates).
-Suppose that PPA holds and that for all sequences $M_{N}\uparrow\infty$ there exist positive constants $C_{1},C_{2},N^{\star}$ such that $\Pi_{r}(\|r-r_{0}\|_{\mathcal{H}}\leq M_{N}N^{-1/2})\geq C_{1}\,e^{-C_{2}\,M_{N}^{2}}$ for $N\geq N^{\star}$.
-Then for all sequences $M_{N}\uparrow\infty$ we have $\mathbb{E}_{\theta_{0}}\Pi_{\alpha}\{H(\theta_{0},\theta)>M_{N}\,N^{-1/2}\mid\boldsymbol{Z}_{N}\}\to 0$ as $N\to\infty$.
+Suppose that PPA holds and that there exist constants $C_{1},C_{2},N^{\star}$ such that $\Pi_{r}(\|r-r_{0}\|_{\mathcal{H}}\leq M_{N}N^{-1/2})\geq C_{1}e^{-C_{2}M_{N}^{2}}$ where $M_{N}/\sqrt{\log N}\to\infty$.
+Then we have $\mathbb{E}_{\theta_{0}}\Pi_{\alpha}\{H(\theta_{0},\theta)>K\,M_{N}\,N^{-1/2}\mid\boldsymbol{Z}_{N}\}\to 0$ as $N\to\infty$ for some sufficiently large $K$.
 It is easy to design priors that satisfy the conditions of Theorem 1.
@@
 Corollary 1 (Rate for Spike-and-Slab).
-Suppose that PPA holds and that $\Pi_{r}(r=r_{0})>0$.
-Then for all sequences $M_{N}\uparrow\infty$ we have $\mathbb{E}_{\theta_{0}}\Pi_{\alpha}\{H(\theta_{0},\theta)\geq M_{N}\,N^{-1/2}\mid\boldsymbol{Z}_{N}\}\to 0$ as $N\to\infty$.
+Suppose that PPA holds, that $\Pi_{r}(r=r_{0})>0$, and that $M_{N}/\sqrt{\log N}\to\infty$.
+Then we have $\mathbb{E}_{\theta_{0}}\Pi_{\alpha}\{H(\theta_{0},\theta)\geq M_{N}\,N^{-1/2}\mid\boldsymbol{Z}_{N}\}\to 0$ as $N\to\infty$.
 Anchoring to the parametric model also enables consistent model selection.
-If the parametric submodel holds and the nonparametric part cannot achieve the parametric rate on its own, then the posterior concentrates all its mass on the parametric model.
+If the parametric submodel holds and the nonparametric part cannot achieve a near-parametric rate on its own, then the posterior concentrates all its mass on the parametric model.
 Theorem 2 (General Model Selection).
@@
 $\Pi^{\star}_{\alpha}(A\mid\boldsymbol{Z}_{N})=\frac{\int_{A}L(\theta)^{\alpha}\ \Pi_{r}^{\star}(dr)\ \Pi_{\eta}(d\eta)}{\int L(\theta)^{\alpha}\ \Pi_{r}^{\star}(dr)\ \Pi_{\eta}(d\eta)},$
-and suppose that $\mathbb{E}_{\theta_{0}}\Pi^{\star}_{\alpha}\{H(\theta_{0},\theta)<M_{N}\,N^{-1/2}\mid\boldsymbol{Z}_{N}\}\to 0$ as $N\to\infty$.
+and suppose that $\mathbb{E}_{\theta_{0}}\Pi^{\star}_{\alpha}\{H(\theta_{0},\theta)<M_{N}\,N^{-1/2}\mid\boldsymbol{Z}_{N}\}\to 0$ where $M_{N}/\sqrt{\log N}\to\infty$ as $N\to\infty$.
 Then $\mathbb{E}_{\theta_{0}}\Pi_{\alpha}(r=r_{0}\mid\boldsymbol{Z}_{N})\to 1$ as $N\to\infty$.
 Remark 1.
-The assumption that $\Pi^{\star}_{\alpha}(\cdot\mid\boldsymbol{Z}_{N})$ fails to achieve the parametric rate is stronger than strictly necessary, as fractional Bayes factors are typically consistent for nested parametric model selection (O’Hagan,, 1995).
-This assumption is not very restrictive, however, because truly nonparametric models rarely achieve parametric rates unless they place positive mass on a correctly specified submodel.
-When Theorem 2 is combined with a parametric submodel $\mathcal{F}_{0}$ that is differentiable in quadratic mean (see Section 5.5 of van der Vaart,, 1998), we also obtain a Bernstein–von Mises result.
+The assumption that $\Pi^{\star}_{\alpha}(\cdot\mid\boldsymbol{Z}_{N})$ fails to achieve a near-parametric rate is stronger than strictly necessary, as fractional Bayes factors are typically consistent for nested parametric model selection (O’Hagan,, 1995) (where the rate is $N^{-1/2}$ for both models).
+We discuss settings where it is obtained for GP priors in Section 3.3, where we note that some GP priors can obtain near-parametric rates; these models fit more naturally with the posterior projection approach described in Section 3.3 in terms of analysis, as a near-parametric rate is useful for proving a Bernstein-von Mises result.
+When Theorem 2 is combined with a parametric submodel $\mathcal{F}_{0}$ that is differentiable in quadratic mean (see Section 5.5 of van der Vaart,, 1998), we also obtain a Bernstein-von Mises result.
 The following theorem formalizes this under standard regularity conditions.
@@
 The value of Theorem 3 relative to the usual Bernstein-von Mises theorem is that it represents an oracle property of the fractional posterior: if the parametric model is correctly specified, our nonparametric prior will nevertheless obtain efficient inference for $\eta_{0}$.
+Efficiency holds up-to inflation of the variance by a factor of $\alpha^{-1}$; this can be corrected using the shift-and-rescale approach of L’Huillier et al., (2023).
 On the other hand, if the parametric model is misspecified we will also be able to obtain efficient nonparametric estimation rates for $\theta_{0}$; the exact rates we get under misspecification will depend on the choice of $\Pi^{\star}(dr)$ and on $r_{0}(x)$.
@@
 Additionally, let $\mathcal{N}$ be the event that all decision trees in the ensemble contain no splitting rules.
-Then $\Pi_{\alpha}(\mathcal{N}\mid\boldsymbol{Z}_{N})\to 1$ in $F_{\theta_{0}}$-probability.
-This result is close in spirit to Theorem 2. While we cannot directly apply Theorem 3, Theorem 4 is strong enough to imply a Bernstein–von Mises result.
+Then $\mathbb{E}_{0}\Pi_{\alpha}(\mathcal{N}\mid\boldsymbol{Z}_{N})\to 1$.
+This result is close in spirit to Theorem 2. While we cannot directly apply Theorem 3, Theorem 4 is strong enough to imply a Bernstein-von Mises result.
 Theorem 5 (BART Bernstein-von Mises).
 Suppose that the conditions of Theorem 4 hold and that $F_{X}$ is such that $X_{i1}\equiv 1$ and that $\mathbb{E}(X_{i}\,X_{i}^{\top})=\Sigma_{X}$ is non-singular.
-Define the vector $\eta=(\beta_{1}+r(0),\beta_{2},\beta_{3},\ldots,\beta_{D},\sigma)^{\top}$ and let $\widehat{\eta}$ by the maximum likelihood estimator of $(\beta,\sigma)$ in Model NR when $r_{0}\equiv 0$.
+Define the vector $\eta=(\beta_{1}+r(0),\beta_{2},\beta_{3},\ldots,\beta_{D},\sigma)^{\top}$ and let $\widehat{\eta}$ be the maximum likelihood estimator of $(\beta,\sigma)$ in Model NR when $r_{0}\equiv 0$.
 Then the conclusion of Theorem 3 holds.
-Remark 2.
-Versions of the above results can also be derived for Model NC and Model DR; we only give results for Model NR for brevity.
-3.3 Results for Gaussian Processes
+3.3 Results for Gaussian Processes and Posterior Projections
 We have discussed two distinct strategies for Gaussian process models.
@@
 For spike-and-slab priors, Corollary 1 applies immediately.
+Remark 2 (Gaussian Process Lower Bounds).
 Applying Theorem 2 directly requires us to have a lower bound on the rate of convergence for Gaussian process models.
-In other words, we need to verify that a pure GP prior without a spike-and-slab component does not achieve the parametric $N^{-1/2}$ rate.
-To keep technical details to a minimum, we find it prudent to adopt $\Pi_{\alpha}\{H(\theta_{0},\theta)>M_{N}\,N^{-1/2}\mid\boldsymbol{Z}_{N}\}\to 0$ in $F_{\theta_{0}}$-probability as a starting assumption rather than to state conditions for it to hold.
-This will almost always be the case when $\kappa(x,x^{\prime})$ is a strictly positive definite kernel such as the squared exponential and Matérn kernels.
-Remark 3 (Gaussian Process Lower Bounds).
-We might expect flexible nonparametric methods to attain a rate of the form $\epsilon_{N}=N^{-p/(2p+D)}$ for some $p>0$.
-For some common kernels (such as the squared exponential kernel), however, we only miss $\epsilon_{N}=N^{-1/2}$ by a logarithmic factor (see Theorem 10 of van der Vaart and van Zanten,, 2011).
-This occurs because the reproducing kernel Hilbert space (RKHS) of the squared exponential kernel consists of extremely smooth functions, making it particularly well suited to approximating the function $r_{0}(x)=0$.
+In other words, we need to verify that a GP prior without a spike-and-slab component does not achieve the near-parametric $(\log N/N)^{1/2}$ rate.
+We might expect flexible nonparametric methods to attain a rate of the form $\epsilon_{N}=N^{-p/(2p+D)}$ up-to logarithmic terms for some $p>0$; this is the case for Matérn priors (see see Theorem 5 of van der Vaart and van Zanten,, 2011 and the discussion thereafter).
+For some common kernels (such as the squared exponential kernel), however, a near-parametric rate is attained (see Theorem 10 of van der Vaart and van Zanten,, 2011).
+This occurs because the squared exponential kernel is supported on extremely smooth functions, making it particularly well suited to approximating the function $r_{0}(x)=0$.
 The general problem of lower bounding posterior contraction rates for Gaussian process regression models is studied by Castillo, (2008).
-If the goal is inference on $\beta_{0}$, it can also suffice to not apply shrinkage and instead project the GP posterior onto the family $\mathcal{F}_{0}$.
+As an alternative to establishing consistent model selection, if the goal is inference on $\beta_{0}$ it can also suffice to instead project the posterior of $(r,\beta)$ onto the family $\mathcal{F}_{0}$.
 Recall that using projections for inference is what we actually recommend.
-For Model NR this is particularly simple due to conjugacy.
+The following two theorems give results for posterior projections of models based on Gaussian process priors; we state both results for the non-fractional posterior.
 Theorem 6 (Semiparametric Bernstein-von Mises Theorem).
@@
 Remark 3.
+Xie and Xu, (2021) study Model NR in detail, and also allow for $x^{\top}\beta$ to be replaced by a nonlinear term $g(x;\beta)$ under appropriate conditions.
+Additionally, they allow $r_{0}(x)\neq 0$ and impose only that $r_{0}(x)$ is sufficiently smooth, with $\Sigma_{X}^{-1}$ replaced with the semiparametric efficient variance in Theorem 6.
+Remark 4.
 For a Gaussian process prior with Model NR, the posterior distribution of $\beta^{\star}$ is normal, and so Theorem 6 really only concerns the behavior of the posterior mean and variance.
 We note that Theorem 6 does not directly rule out that a Bernstein-von Mises result holds without the linear term $\sigma^{2}_{\beta}x^{\top}x^{\prime}$ in the kernel.
-In Section 4.3 we confirm empirically that Theorem 6 accurately predicts performance, while kernels without the linear term are empirically very highly biased in finite samples.
-Theorem 6 is proven using the conjugacy of the GP prior.
-To show that the result is general, we also prove an analogous result for logistic regression (Model NC).
+In Section 4.3 we confirm empirically that Theorem 6 accurately predicts performance, while kernels without the linear term are empirically highly biased in finite samples.
+To show that this phenomenon is general, we prove an analogous result for logistic regression (Model NC).
 The proof follows Theorem 2.2 of L’Huillier et al., (2023), which is itself based on Theorem 2.1 of Castillo and Rousseau, (2015).
-The main ingredients of the proof are that (i) $\|r-r_{0}\|=o(N^{-1/4})$ with high posterior probability, which allows us to control the remainder of a Taylor expansion and (ii) $r_{0}(x)=x^{\top}\beta_{0}$ lies in the RKHS of the prior because of the linear term in the kernel.
-Theorem 7 (Logistic GP Bernstein–von Mises).
-Consider Model NC with $r_{0}(x)\equiv 0$ and $X_{1},X_{2},\ldots$ a deterministic sequence in $0,1]^{P}$.
-Suppose $\widehat{I}_{N}=N^{-1}\sum_{i}\mu_{0}(X_{i})\{1-\mu_{0}(X_{i})\}X_{i}X_{i}^{\top}$ is invertible and $\widehat{I}_{N}\to I$ for some positive-definite $P\times P$ matrix $I$.
-Let $\Pi(\cdot\mid\boldsymbol{Z}_{N})$ be the posterior distribution of $r(\cdot)$ associate to the model with $Y_{i}\mid X_{i},r]\sim\operatorname{Bernoulli}\{(1+e^{-r(X_{i})})^{-1}\}$ and $r\sim\operatorname{GP}(0,\kappa)$ where $\kappa(x,x^{\prime})=\sigma_{\beta}^{2}x^{\top}x^{\prime}+\kappa^{\star}(x,x^{\prime})$ for $\sigma_{\beta}^{2}>0$ and $\kappa^{\star}$ satisfying Assumption A in the Supplementary Material.
-Define $\beta^{\star}=\widehat{I}_{N}^{-1}N^{-1}\sum_{i}\mu_{0}(X_{i})\{1-\mu_{0}(X_{i})\}X_{i}r(X_{i})$.
+The main steps of the proof are (i) we assume that $\|r-r_{0}\|_{N}=\sqrt{N^{-1}\sum_{i}\mu_{0}(X_{i})\{1-\mu_{0}(X_{i})\}\{r_{0}(X_{i})-r(X_{i})\}^{2}}$ has a posterior contraction rate faster than $N^{-1/4}$, which allows us to control the remainder of a Taylor expansion and (ii) a change-of-variables argument under the prior $\beta\sim\operatorname{Normal}(\mu_{\beta},\Sigma_{\beta})$, similar to the application of the Cameron-Martin theorem by Castillo and Rousseau, (2015).
+On the technical side, guaranteeing that item (i) holds is difficult due to the fact that $\|r-r_{0}\|_{N}$ is not equivalent to the empirical Hellinger distance, and so is not available from existing results for Gaussian process logistic regression; as discussed in the Supplementary Material, this difficulty can be bypassed by working with an appropriately-truncated prior instead.
+As with the results of Xie and Xu, (2021), we expect that this result can be modified with additional assumptions to also accommodate $r_{0}(x)$ deviating from $x^{\top}\beta_{0}$ provided that $\|r-r_{0}\|_{N}$ has a posterior contraction rate faster than $N^{-1/4}$; after such a modification, the asymptotic variance would be the semiparametric efficient variance.
+Theorem 7 (Bernstein-von Mises for Logistic Regression).
+Consider Model NC with $r_{0}(x)\equiv 0$ and consider a fixed sequence $X_{1},X_{2},\ldots$ in $0,1]^{P}$ and let $\widehat{I}_{N}=\frac{1}{N}\sum_{i=1}^{N}\mu_{0}(X_{i})\{1-\mu_{0}(X_{i})\}X_{i}\,X_{i}^{\top}$.
+Suppose $\widehat{I}_{N}$ is invertible and that $\widehat{I}_{N}\to I$ for some positive definite matrix $I$.
+Let $\Pi(\cdot\mid\boldsymbol{Z}_{N})$ be the posterior distribution of $(r,\beta)$ associated to the model with $Y_{i}\mid X_{i},r,\beta]\sim\operatorname{Bernoulli}\{(1+e^{-(X_{i}^{\top}\beta+r(X_{i}))})^{-1}\}$ such that Assumption A in the Supplementary Material is satisfied.
+Define $\beta^{\star}=\widehat{I}_{N}^{-1}N^{-1}\sum_{i}\mu_{0}(X_{i})\{1-\mu_{0}(X_{i})\}X_{i}\{r(X_{i})+X_{i}^{\top}\beta\}$ where $\widehat{I}_{N}=N^{-1}\sum_{i}\mu_{0}(X_{i})\{1-\mu_{0}(X_{i})\}X_{i}\,X_{i}^{\top}$.
 Then
-$d\left\{\sqrt{N}(\beta^{\star}-\widehat{\beta}),\operatorname{Normal}(0,I^{-1})\right\}\to 0\quad\text{in $F_{\theta_{0}}$-probability}$
+$d\left\{\Pi\left(\sqrt{N}(\beta^{\star}-\widehat{\beta})\in\cdot\mid\boldsymbol{Z}_{N}\right),\operatorname{Normal}(0,I^{-1})\right\}\to 0\quad\text{in $F_{\theta_{0}}$-probability}$
 where $d(\cdot,\cdot)$ metrizes convergence in distribution and $\widehat{\beta}=\beta_{0}+N^{-1}\widehat{I}_{N}^{-1}\sum_{i=1}^{N}\{Y_{i}-\mu_{0}(X_{i})\}X_{i}$.
-Theorem 7 shows that a certain “information-weighted” projection $\beta^{\star}$ satisfies a semiparametric Bernstein–von Mises theorem with $I$ the efficient Fisher information under $\mathcal{F}_{0}$.
-The quantity $\widehat{\beta}$ is equivalent to the MLE under the parametric logistic regression submodel, being a one-step Newton–Raphson update of a $\sqrt{N}$-consistent quantity.
-We use the information-weighted projection mainly for technical convenience, since it makes $\beta^{\star}$ a linear functional of $r(\cdot)$.
-In practice, it is much better to use the Kullback-Leibler projection $\beta^{\star}=\arg\min_{\beta}\sum_{i}\mu(X_{i})\log\mu(X_{i})/\mu_{\beta}(X_{i})$ where $\mu_{\beta}(x)=(1+e^{-x^{\top}\beta})^{-1}$, as in experiments we have found it to be far less biased.
+Theorem 7 shows that a certain “information-weighted” projection $\beta^{\star}$ satisfies a semiparametric Bernstein-von Mises theorem with $I$ the efficient Fisher information under $\mathcal{F}_{0}$.
+The quantity $\widehat{\beta}$ is asymptotically equivalent to the MLE under the parametric logistic regression submodel, being a one-step Newton–Raphson update of a $\sqrt{N}$-consistent quantity.
+We use the information-weighted projection mainly for technical convenience, as it makes $\beta^{\star}$ a linear functional of $\theta(x)=r(x)+x^{\top}\beta$.
+In practice, it is much better to use the Kullback-Leibler projection $\beta^{\star}=\arg\min_{\beta}\sum_{i}\mu(X_{i})\log\frac{\mu(X_{i})}{\mu_{\beta}(X_{i})}+\{1-\mu(X_{i})\}\log\frac{1-\mu(X_{i})}{1-\mu_{\beta}(X_{i})}$ where $\mu_{\beta}(x)=(1+e^{-x^{\top}\beta})^{-1}$, as in experiments we have found it to be far less biased.
 4 Applications
@@
 We see that when $\lambda_{0}=0$, the performance of GBART is indistinguishable from that of the linear model at all sample sizes.
-Somewhat surprisingly, for smaller signal-to-noise ratios and sample sizes the unmodified BART model outperforms a correctly-specified linear model; this is because the sample size is not large enough to estimate the coefficients reliably, giving an advantage to models that regularize the predictions towards zero.
+Somewhat surprisingly, for smaller signal-to-noise ratios and sample sizes the unmodified BART model outperforms a correctly-specified linear model; this is because the sample size is not large enough to estimate the coefficients reliably, giving an advantage to models that regularize the predictions toward zero.
 This defect in the linear model can be mitigated by replacing the least squares estimates with a ridge regression or lasso.
 When $\lambda_{0}=0.4$, GBART outperforms the other methods, particularly at large sample sizes.
-It is strictly superior to the linear model under all data generating mechanisms.
+Under these simulation settings, it is strictly superior to the linear model under all data generating mechanisms.
 We again see that when both the sample size and the signal-to-noise ratios are small there is an advantage to using a BART model that more strongly regularizes the predictions.
@@
 When $\lambda_{0}=0$ the parametric model is consistently selected, while as $\lambda_{0}$ increases the nonparametric model is selected with high probability.
-The trend is more muted when $\sigma_{0}$ is larger; this is intuitive, since (for example) $\sigma_{0}=4$ corresponds to a signal-to-noise ratio of roughly 16-to-1, while $\sigma_{0}=1$ corresponds to a signal-to-noise ratio of roughly 1-to-1.
+The trend is more muted when $\sigma_{0}$ is larger; this is intuitive, since (for example) $\sigma_{0}=4$ corresponds to a signal-to-noise ratio of roughly 1-to-16, while $\sigma_{0}=1$ corresponds to a signal-to-noise ratio of roughly 1-to-1.
 As a general lesson, these results suggest that projection-based inference is often preferable to relying solely on model selection consistency.
@@
 We examine this in more detail in Section 4.3.
-4.3 Semiparametric Bernstein–von Mises for Gaussian Processes
-We now empirically validate the Bernstein–von Mises result for the projection of the GP posterior onto the linear regression model. We use the data generating process
+4.3 Semiparametric Bernstein-von Mises for Gaussian Processes
+We now empirically validate the Bernstein-von Mises result for the projection of the GP posterior onto the linear regression model. We use the data generating process
 $Y_{i}=\sum_{j=1}^{P}\beta_{0j}X_{ij}+\epsilon_{i},$
@@
 We then fit a GP regression model $Y_{i}=\mu(X_{i})+\epsilon_{i}$ with $\mu\sim\operatorname{GP}(0,\kappa)$ and compute the posterior distribution of the projection parameter $\beta^{\star}=(\boldsymbol{X}^{\top}\boldsymbol{X})^{-1}\boldsymbol{X}^{\top}\boldsymbol{\mu}$.
-We consider the Laplace kernel $\kappa(x,x^{\prime})=e^{-\|x-x^{\prime}\|}$, the squared exponential kernel $\kappa(x,x^{\prime})=e^{-\|x-x^{\prime}\|^{2}}$, and the squared-exponential-plus-linear kernel $\kappa(x,x^{\prime})=100\,(x^{\prime})^{\top}x+e^{-\|x-x^{\prime}\|_{2}}$.
+We consider the Laplace kernel $\kappa(x,x^{\prime})=e^{-\|x-x^{\prime}\|}$, the squared exponential kernel $\kappa(x,x^{\prime})=e^{-\|x-x^{\prime}\|^{2}}$, and the squared-exponential-plus-linear kernel $\kappa(x,x^{\prime})=100\,(x^{\prime})^{\top}x+e^{-\|x-x^{\prime}\|_{2}^{2}}$.
 Figure 4: Results for the experiments in Section 4.3.
@@
 We monitor three quantities: (i) the nominal coverage of 95% credible intervals for $\beta^{\star}_{1}$ under each kernel; (ii) the scaled bias of the posterior mean, defined as the average of $\sqrt{N}(\widehat{\beta}_{1}-\beta_{01})$ with $\widehat{\beta}=\mathbb{E}_{\Pi}(\beta^{\star}\mid\boldsymbol{Z}_{N})$; and (iii) the scaled variance, defined as the average of $N\operatorname{Var}_{\Pi}(\beta^{\star}_{1}\mid\boldsymbol{Z}_{N})$.
-If the posterior satisfies a semiparametric Bernstein–von Mises result, we expect coverage near 95%, scaled bias near zero, and scaled variance near 1, since $\mathbb{E}_{\theta_{0}}(\boldsymbol{X}^{\top}\boldsymbol{X})=\mathrm{I}$ in this setup.
+If the posterior satisfies a semiparametric Bernstein-von Mises result, we expect coverage near 95%, scaled bias near zero, and scaled variance near 1, since $\mathbb{E}_{\theta_{0}}(N^{-1}\boldsymbol{X}^{\top}\boldsymbol{X})=\mathrm{I}$ in this setup.
 Results
 Figure 4 shows results across kernels and sample sizes.
-Among the three, only the squared-exponential-plus-linear kernel empirically satisfies the semiparametric Bernstein–von Mises theorem at the chosen sample sizes.
+Among the three, only the squared-exponential-plus-linear kernel empirically satisfies the semiparametric Bernstein-von Mises theorem at the chosen sample sizes.
 In larger samples, we do see the asymptotic variance of the squared exponential and Laplace kernels approaching the semiparametric efficiency, and indeed asymptotically we do attain the semiparametric efficiency level; the problem with these kernels is that, even in very large samples, the estimates are highly biased, which is clearly seen in the middle panel of Figure 4.
@@
 We now show how to diagnose model misspecification in practice by analyzing data from the Medical Expenditure Panel Survey (MEPS, Cohen et al.,, 2009).
-We consider a subset of the data from the 2011 survey containing 16,113 with the goal of predicting self-assessed health status.
+We consider a subset of the 2011 survey with 16,113 observations, with the goal of predicting self-assessed health status.
 We use Model NR with $Y_{i}$ representing the self-assessed health status of individual $i$ and $X_{i}$ representing the individual’s age, body mass index (BMI), education level, income relative to the poverty line, region of the country, sex, marital status, race, seatbelt use, and smoking status.
-To illustrate the ability of the GBART model to default to a linear model when there is little data while also adapting to non-linearities and interactions in the data, we consider two analyses: one using a randomly selected sample of size of $N=1000$ and one using a sample of size $N=8000$.
+Although self-assessed health status is ordinal, here we use Model NR as a simple working Gaussian model for prediction and projection-based summarization; extending the same workflow to ordinal or logistic-style models is also possible.
+To illustrate the ability of the GBART model to default to a linear model when there is little data while also adapting to nonlinearities and interactions in the data, we consider two analyses: one using a randomly selected sample of size $N=1000$ and one using a sample of size $N=8000$.
 For both cases we monitor the summary $R^{2}$ of the projection onto a linear regression.
+All projection summaries reported below are in-sample summaries computed on the analyzed sample.
 To assess how closely the semiparametric Bernstein-von Mises result holds, we also monitor the posterior distribution of the coefficient of BMI.
@@
 Results for the $N=1000$ model are given in Figure 5.
-We see that the model is shrunk very heavily towards a linear regression model, with the posterior summary $R^{2}$ being very close to $100\%$.
+We see that the model is shrunk very heavily toward a linear regression model, with the posterior summary $R^{2}$ being very close to $100\%$.
 The posterior distribution of the coefficient of BMI is also very close to the posterior from a parametric Bayesian linear model, indicating that the semiparametric Bernstein-von Mises approximation is accurate.
@@
 Figure 7: For the $N=8000$ additive model: the posterior summary $R^{2}$ and overall summary $R^{2}$ of the linear model (left) and plot of the projected predictions against the actual predictions from the model (right).
-From the analysis for the $N=8000$ setting, we see that there is evidence of misspecification of the linear regression, implying that there exist either non-linearities or interactions in the data.
+From the analysis for the $N=8000$ setting, we see that there is evidence of misspecification of the linear regression, implying that there are either nonlinearities or interactions in the data.
 An advantage of the fully-Bayesian approach is that we are free to summarize the posterior in any way we choose, and in particular we can look for better models to project the posterior of $\mu(x)$ onto.
@@
 After making this modification, we find that the overall summary $R^{2}$ of the model exceeds 98%, with the posterior distribution of $R^{2}$ centered around 96%, giving a roughly $10\%$ increase in predictive variability accounted for.
-Figure 8: Decision tree fit to the residuals $\widehat{\mu}(X_{0})-\widetilde{\mu}(X_{i})$ from the additive model.
+Figure 8: Decision tree fit to the residuals $\widehat{\mu}(X_{i})-\widetilde{\mu}(X_{i})$ from the additive model.
 To gain better insight into the remaining sources of variability, Figure 8 displays a CART fit to the residuals $\widehat{\mu}(X_{i})-\widetilde{\mu}(X_{i})$, where $\widetilde{\mu}(x)$ denotes the additive summary of $\widehat{\mu}(x)$.
-The magnitude of the predictions of the CART model represent how off, on average, the additive summary is from the general BART model for subsets of the observations.
-The largest degree of misfit occurs for individuals who are white and with poverty level less than $121$, with the additive summary predicting a health status on average 0.036 units higher than the GBART model.
+The CART predictions quantify the average discrepancy between the additive summary and the GBART fit within each subgroup.
+The largest degree of misfit occurs for white individuals with poverty level below $121$, with the additive summary predicting a health status on average 0.036 units higher than the GBART model.
 Summary
@@
 5 Discussion
-The position advocated for in this manuscript is one of defensive model expansion, wherein we embed a parametric Bayesian model inside a nonparametric Bayesian model and apply strong shrinkage towards the parametric model.
+The position advocated for in this manuscript is one of defensive model expansion, wherein we embed a parametric Bayesian model inside a nonparametric Bayesian model and apply strong shrinkage toward the parametric model.
 We claim that this retains many of the benefits of specifying a parametric Bayesian model, such as strong performance when $N$ is small or the signal-to-noise ratio is low, while mitigating the drawbacks of potential model misspecification.